### v3

# Review of Demographic Fairness in Face Recognition

Abstract
The issue of difference in face recognition (FR) performance across demographic groups has emerged as a critical area of research, given its impact on fairness, equity, and reliability across diverse applications. As FR technologies are increasingly deployed globally, disparities in performance across demographic groups– such as race, ethnicity, and gender– have garnered significant attention. These differences or biases not only compromise the credibility of FR systems but also raise ethical concerns, especially when these technologies are employed in sensitive domains. This review consolidates extensive research efforts providing a comprehensive overview of the multifaceted aspects of demographic fairness in FR.
We systematically examine the primary causes, datasets, assessment metrics, and mitigation approaches associated with performance differences in FR across demographic groups. By categorizing key contributions in these areas, this work provides a structured approach to understanding and addressing the complexity of this issue. Finally, we highlight current advancements and identify emerging challenges that need further investigation. This article aims to provide researchers with a unified perspective on the state-of-the-art while emphasizing the critical need for equitable and trustworthy FR systems.
Index Terms:
Demographic Fairness, Bias, Face Recognition, Biometrics, Differential Performance, Trustworthy AI
I Introduction
Differences in the performance of face recognition (FR) system across demographic groups, commonly described under the umbrella of demographic fairness, have emerged as a critical challenge in the deployment of biometric technologies for real-world applications [1, 2, 3, 4]. Several real-world incidents underscore the societal risks associated with such disparities. For example, a Black individual was wrongly arrested in Detroit due to a false FR match [5]; a US civil liberties test reported misidentification of a handful US lawmakers as criminals by major FR service [6]; and recently, researchers exposed racial bias in airport FR systems used during passenger boarding [7, 8]. These systems frequently exhibit unequal recognition performance across race, gender, and age groups [4, 9, 10]. Such disparities, often labelled as biases, can have far-reaching consequences, especially in critical applications like border crossing, law enforcement [11, 12], security [13], and hiring processes [14, 15, 16], where both fairness and accuracy are paramount.
Figure 1: Illustration of differences in performance across demographic groups in FR with reference to different demographic factors: (a) race or ethnicity, (b) gender, and (c) age. Score distributions for mated and non-mated pairs are shown for multiple demographic groups, along with decision thresholds (dotted lines) corresponding to different operating points. This figure is intended to demonstrate that accuracy differentials may persist across a range of thresholds and are not restricted to a particular type of error (e.g., false match or false non-match).
A conventional FR system evaluates the similarity between two facial images by computing a matching score based on their extracted features to determine whether they correspond to the same individual or not. The performance of such systems is typically examined by analyzing the distributions of matching scores for image pairs of the same identity and those of different identities. Scores are thresholded into binarized decisions: match or no-match. Fig. 1 illustrates the issue of demographic bias based on race, gender, and age. Ideally, a fair FR model should exhibit equitable performance across all demographic groups (i.e., similar error rates or comparable score distributions). A commonly implicit but rarely emphasized assumption is that the groups being compared are equivalent in all aspects except for the demographic attribute of interest. For instance, comparing a group of persons wearing face masks with a group without masks- one may introduce additional variability unrelated to demographics, making the comparison unsuitable for assessing fairness. The disparity between equivalent groups contributes to the unfair or biased FR. (The formal definitions for error rates and disparities provided in Sec. II and V.) As depicted by various choices of scoring thresholds, accuracy differences exist across demographic splits and can manifest at various thresholds. This emphasizes that demographic disparities are not tied solely to false match or false non-match errors but may be present throughout the score space. This issue is further compounded by the growing reliance on FR technologies, making it imperative to identify, evaluate, and mitigate sources of accuracy differences effectively.
Due to its severity and the widespread range of FR applications, demographic fairness has emerged as a crucial area of research, drawing significant attention from both the biometrics and computer vision communities [3, 4, 17, 18]. This issue has been formally incorporated into the evaluation frameworks of prominent initiatives. The National Institute of Standards and Technology (NIST) is a US agency that independently evaluates biometric technologies. Its Face Recognition Vendor Tests (FRVT) benchmark the accuracy of FR algorithms using standardized procedures and extensive datasets. Since 2019, FRVT reports have incorporated analyses of demographic disparities [19, 20], making them a key reference for assessing fairness in FR. Other initiatives, including the Maryland Test Facility (MdTF) and the European Association for Biometrics (EAB), have also contributed to this discourse, though at a smaller scale and scope compared to NIST. The MdTF, supported by the United States Department of Homeland Security (DHS), has conducted biometric technology rallies to evaluate demographic disparities in FR systems. In Europe, organizations such as the EAB have hosted dedicated events on demographic fairness in biometric systems, underlining the global importance of this topic. This research area is often positioned within the broader context of fair and trustworthy biometrics, and has been receiving substantial attention– in the form of papers, workshops, or special sessions– from leading conferences such as IEEE/CVF CVPR, WACV, IEEE FG, ICPR, ICLR, and ICML; and from reputable journals including IEEE Transactions on Information Forensics and Security (TIFS), IEEE Transactions on Biometrics, Behavior, and Identity Science (TBIOM) [21] and IEEE Signal Processing Magazine [22]. Additionally, standards organizations have recognized the need for systematic approaches. For example, the recently published ISO/IEC 19795-10 guidelines quantify demographic differentials in biometric systems, and emphasize the need for addressing fairness in FR technologies [23].
Scope: Fairness and bias in machine learning are expansive topics, and their application in biometrics has drawn significant attention in recent years. Several comprehensive reviews have addressed fairness and bias in machine learning broadly [9, 24], while others focus specifically on biometrics, offering insights into various modalities such as face, fingerprint, and vein, alongside applications beyond recognition, including region of interest (ROI) detection, quality assessment, and presentation attack detection[3, 4, 17, 25]. However, as face remains the most commonly used biometric trait in operational scenarios, a substantial portion of research on differences across demographic performance or bias has concentrated on this modality.
While previous reviews offer broad perspectives, the extensive literature and emerging challenges specific to demographic fairness in FR necessitate a dedicated review. In this article, we provide a consolidated discussion of recent advancements in the field, addressing the causes of demographic bias, available datasets for research, evaluation metrics for assessment of fairness, and recent mitigation techniques. Furthermore, we explore ongoing challenges that persist in addressing demographic disparities, particularly in the light of novel use cases and emerging FR applications. Although our primary focus is on race and ethnicity, as these are dominant areas of research, we also include gender-related studies within the broader context. Age-related studies, given their distinct nature (variation over time for an individual) and established body of research, are referenced only where directly relevant.
Naming Conventions: The primary objective of this review is to examine performance differences in FR across demographic groups. Here, the term performance is used broadly to encompass any metric relevant to FR evaluation, although it commonly refers to accuracy. The term differences denotes a quantifiable amount of the given performance metric measured on a given dataset. In some instances, we use the term demographic bias to describe performance differences observed across demographic groups (e.g., race, gender, age), which may be reflected in metrics such as accuracy or error rates. This terminology is used to align with common usage in recent literature and is not meant to suggest deliberate discrimination or convey any moral judgment.
Additionally, terms such as race and ethnicity are often used interchangeably, although they represent distinct concepts. For clarity, we retain the terms employed by the original studies. Similarly, the names of ethnic groups vary across the literature: for instance, some works use terms like Black and White, while others prefer African (or African-American) and Caucasian. Additionally, the South Asian group is sometimes referred to as Indian, whereas Asians often refer to East Asians. To maintain consistency and respect the source material, we adhere to the original terminologies in this review. Readers are encouraged to refer to the cited works for precise definitions and context. While fairness in related topics such as face detection, image quality, expression recognition, and attribute estimation is of interest to the research community, this review exclusively focuses on demographic fairness in FR.
Contributions: This review constitutes a comprehensive work dedicated to exploring demographic fairness in FR, offering a unified and holistic perspective to researchers in the field. We systematically analyze and organize key aspects, including the causes of demographic disparities, available datasets, assessment metrics, and mitigation techniques, providing a structured framework for understanding these areas. Finally, we identify emerging challenges and unresolved questions, inviting further research and innovation to advance equitability in FR systems.
The structure of this paper is as follows: We begin with preliminaries of assessment of generic FR in Section II. Section III explores the causes of differences in performance, analyzing factors such as distribution of demographic groups in datasets, skin-tone, image quality, and soft attributes. We provide an overview of datasets commonly used for fairness-related research in Section IV, highlighting their demographic attributes and suitability for specific tasks. Section V reviews existing metrics for fairness evaluation, discussing their strengths and limitations. Section VI outlines recent bias mitigation strategies across different stages of the FR pipeline. We discuss open challenges and future research directions in Section VII, and conclude the review in Section VIII.
II Preliminaries of FR Evaluation
Before delving into the fairness aspect of FR systems, we briefly outline the standard FR pipeline and its associated evaluation terms. In a typical verification setting, two face images are first processed through detection and alignment stages to conform to the FR model’s input requirements. The aligned faces are then passed through a feature extractor, usually a deep neural network, which produces a compact embedding–a low-dimensional representation of the subject’s facial features. The similarity between two such embeddings is computed (commonly using cosine similarity or Euclidean distance), resulting in a continuous score typically bounded in [0, 1]. In many applications, particularly those involving enrolled users (e.g., access control), the embeddings are precomputed and stored in a secure gallery. For identification tasks, the query embedding is matched against a set of gallery templates to determine the identity.
In this context, a genuine match (or mated pair) refers to two samples originating from the same individual, while an impostor match (or non-mated pair) involves samples from different individuals. The similarity score is binarized using a threshold (often denoted as $\tau$), producing a match or non-match decision. The False Match Rate (FMR) is the proportion of impostor pairs incorrectly classified as matches (false positives), while the False Non-Match Rate (FNMR) is the proportion of genuine pairs incorrectly classified as non-matches (false negatives). These two metrics are fundamental to evaluating the accuracy and reliability of FR systems. Their values depend on the chosen threshold $\tau$, with lower FMR and FNMR generally indicating better performance. The overlap between the genuine and impostor score distributions (see Fig. 2) defines the regions where these classification errors arise.
Figure 2: Illustration of false matches and false non-matches arising from distributions of mated and non-mated scores along with the score threshold.
The selection of the score threshold has a direct impact on the error rates in face recognition. Increasing the threshold tends to reduce the False Match Rate (FMR) by making it harder for impostor pairs to be incorrectly accepted as matches, but this comes at the cost of a higher False Non-Match Rate (FNMR), as more genuine pairs may be rejected, and vice versa. It is important, however, to distinguish between the inherent overlap in the score distributions and the binary classification decision that results from thresholding. The overlap in the range of similarity scores between mated and non-mated pairs indicates intrinsic ambiguity in the matching process, while the threshold simply determines the point of separation between match and non-match decisions.
Two demographic groups may exhibit a similar degree of score overlap yet yield different FMR and FNMR values due to differences in the shapes of their score distributions or disparities in group-specific population characteristics. Conversely, distributions across groups may differ significantly, while their overall error rates remain similar. These variations– in either score distribution overlap or resulting error rates– across demographic subgroups form the foundation of what is commonly referred to as bias in FR systems. This review focuses on analyzing, evaluating, and mitigating these disparities.
III Causes of Performance Differences
In this section, we consolidate findings from existing works related to causes of performance differences in FR. Considering the wide range of research in this area, we have grouped the causes into categories for clarity and systematic review. These categories encompass factors such as imbalances in training datasets, variability in skin-tones, algorithmic sensitivity, image quality, related covariates, combined or intersectional factors, and soft attributes. Though this categorization simplifies the organization, it is important to note that many studies attribute accuracy differences to multiple, overlapping factors, making strict classification difficult. We have categorized works based on their primary focus or findings, but some studies have been referenced across multiple categories to reflect their broader relevance. This approach also highlights that causes of demographic differences or bias are inherently multifaceted and interconnected, requiring a thorough understanding to analyze and mitigate these disparities. A quick summary of various works discussed in this section can be found in Table I.
III-A Training Datasets
Early studies in this area speculated that the issue of performance differences in FR systems across demographic groups often stems from imbalanced or unrepresentative datasets, significantly influencing both training and evaluation outcomes. Research by Krishnapriya et al. [26] demonstrated how demographic groups, such as African-American cohorts, exhibit higher FMRs, while Caucasian cohorts face higher FNMR, highlighting the interplay between race and matching thresholds. The FRVT conducted by NIST [19] reported increased false positive identifications in women, children, and the elderly, alongside higher false negatives in under-represented racial groups– particularly Black, East Asian, and Native American individuals– emphasizing the intricate interactions between dataset characteristics and demographic attributes.
In a pre-deep learning era, studies like Klare et al.  [27] examined the impact of demographic balance in training datasets on FR performance. While their results indicated that underrepresented groups– such as females and Black individuals– benefited from increased representation, they also found that balancing the dataset alone did not fully eliminate disparities in recognition accuracy at test time. To further mitigate this, they proposed training exclusive cohort-specific models and using dynamic model selection during inference, highlighting that data balance might often be complemented with architectural or procedural adjustments to meaningfully reduce accuracy differences. Cavazos et al. [28] identified how dataset complexity and identification thresholds contribute to racial bias, such as the need for higher thresholds for East-Asian faces to achieve comparable false acceptance rates (FARs). Gwilliam et al. [29] challenged the prevailing assumptions about the necessity of balanced datasets by demonstrating that skewed distributions favoring African faces reduced racial differentials more effectively than balanced datasets. In [30, 31], Albiero et al. demonstrated that balancing training data for both genders (male and female) did not help reducing the corresponding differential in FR accuracy. Wu and Bowyer [32] expanded this discussion, emphasizing that mere balance in identities or number of images is insufficient to address disparities, highlighting additional factors like brightness and head pose during dataset assembly. These factors–often referred to as soft attributes–include physical traits, behavioral patterns, or material accessories associated with a person. They are descriptive and can be categorized or classified [33]. The soft-biometric attributes tend to influence the performance of FR systems.
Other works delved into specific aspects of demographic balance in datasets. Wang et al.  [34] observed that even race-balanced datasets failed to eliminate accuracy differentials, hypothesizing that certain ethnicities are inherently more challenging to recognize. Kolla and Savadamuthu [35] highlighted the influence of facial quality and racial feature gradations on model fairness. Focusing on inter-sectional bias, Muthukumar et al.  [36] identified structural facial features as significant contributors, particularly for dark-skinned females, over attributes like skin-tone or hair length. Cook et al. [37] further analyzed the role of image acquisition conditions, noting how factors such as skin reflectance and environmental conditions disproportionately affect darker-skinned individuals, thus advocating for standardized acquisition protocols to mitigate differences in FR performance.
Although several studies acknowledge that significantly imbalanced training data contribute to unfair or biased FR models and training with balanced datasets reduces disparities to some extent, there is a consensus that these are neither the sole causes nor complete solutions to the broader issue of demographic fairness. Availability of large-scale training data, coupled with deeper models with better learning capacity, has made it possible to investigate the aspect of datasets in a more detailed manner. Recent findings have indicated that the role of balancedness of training data is limited, and other factors that are inherent to the individuals or their soft attributes may have a bigger role. Additionally, balancing training data (especially through sub-sampling majority groups) may involve discarding large amounts of high-quality data, thereby reducing overall training volume and potentially decreasing recognition accuracy while reducing the accuracy across constituent demographic groups.
III-B Variability in Skin-tone
The influence of skin-tone on the performance of FR systems has been extensively studied, revealing significant demographic disparities. A common method to categorize skin-tone for FR studies is the Fitzpatrick Skin Type Classification, which divides skin types into 6 categories (Type I to VI) based on skin’s reaction to ultraviolet (UV) exposure and general pigmentation [38]. While widely used due to its simplicity, it was originally designed for dermatology, not computer vision. Alternative categorization schemes include the von Luschan’s chromatic scale, a 36-point system based on observed skin color using standardized tiles, and the Monk Skin-Tone (MST) Scale, a 10-tone scale designed to represent a broader and more inclusive range of skin-tones [39]. Each scheme varies in granularity and applicability, and ongoing research continues to assess their effectiveness.
In [36], Muthukumar et al. identified notable under-performance in recognizing dark-skinned females compared to other demographic group for commercial classifiers. Their analysis attributed these disparities to structural features such as lips, eyes, and cheeks, in addition to skin-tone itself. In another study, Krishnapriya et al.  [26] analyzed FMR and FNMR across skin-tone groups and observed error rate variations among demographic cohorts; however, they did not find conclusive evidence linking darker skin-tones directly to higher error rates. Buolamwini and Gebru [40] employed the Fitzpatrick skin classification system to evaluate commercial gender classifiers and reported the lowest accuracy for darker-skinned females. While this study evaluated gender classification rather than identity recognition, it led to broader discussions around fairness in facial analysis systems, including those used in biometric applications.
The Biometric Technology Rallies organized by MdTF have offered comprehensive insights into the role of skin-related factors impacting FR performance. Their 2019 report [37] emphasized skin reflectance as a more significant predictor of performance disparities than race. Skin reflectance refers to the measurable amount of light reflected from the skin surface, which affects how well facial features are captured in images; whereas skin-tone refers to perceived skin color [37]. Using systematic linear modeling, their study demonstrated that darker skin-tones were associated with longer transaction (processing overall pipeline) times and lower accuracy in biometric systems. The longer transaction times were primarily attributed to difficulties in the face detection or image acquisition stage under suboptimal lighting conditions. Lower skin reflectance can reduce contrast, making it harder for detection algorithms to localize the face, thus increasing processing time. This dependency was found to vary substantially across systems, highlighting important role of acquisition methods in determining the extent of performance differences. Lu et al.  [41] provided a quantitative assessment of performance variations across five skin-tone groups, identifying light-skinned individuals as the easiest to verify and darker-skinned individuals as the most challenging. However, ambiguities in defining skin-tone categories complicate direct evaluations, highlighting the need for standardized classification metrics.
While many studies report that individuals with lighter skin-tones tend to be recognized more accurately than those with darker skin-tones, there is no consistent consensus that skin-tone is the primary driver of differences in FR performance across demographic groups. Moreover, it is important to note that most analyses rely on apparent or perceived skin tone–estimated via brightness or proxy classification schemes–rather than direct, objective measurement.
TABLE I: Summary of works delving into various causes of demographic fairness in face recognition. As several works have identified multiple causes of such performance differences, we have categorized the works based on their primary focus or inference. For details, readers are encouraged to refer to the source materials. The demographic factors of primary interest are denoted as ET: Ethnicity or race, GN: gender or sex, AG: age; whereas + indicates study of more attributes.
Reference Year Dataset Attribute Summary
Training Datasets Klare et al. [27] 2012 PCSO ET, GN, AG Advocated for balanced datasets and exclusive cohorts to improve face recognition performance.
Muthukumar et al. [36] 2018 PPB GN, ST Identified structural facial features as primary contributors to intersectional bias for dark-skinned females.
NIST FRVT [19] 2019 Private ET, GN, AG, + Reported increased false positives in women, children, and elderly, and higher false negatives in underrepresented groups.
Wang et al. [34] 2019 RFW ET Observed that race-balanced datasets do not fully eliminate bias, suggesting inherent challenges in recognizing certain ethnicities.
Cook et al. [37] 2019 Private ET, GN, AG, + Analyzed image acquisition conditions, noting the impact of skin reflectance and environmental factors on darker-skinned individuals.
Krishnapriya et al. [26] 2020 MORPH ET, ST Demonstrated demographic disparities in FMR and FNMR, with African-American cohorts having higher FMR and Caucasian cohorts higher FNMR.
Cavazos et al. [28] 2020 GBU ET Highlighted how dataset complexity and thresholds affect racial bias, requiring higher thresholds for East-Asian faces.
Gwilliam et al. [29] 2021 BUPT, RFW ET Showed that skewed distributions favoring African faces can mitigate racial bias better than balanced datasets.
Wu and Bowyer [32] 2023 DemogPairs, RFW, BFW, ET, GN, + Emphasized that balancing identities and images alone is insufficient, stressing brightness and head pose considerations.
Kolla and Savadamuthu [35] 2023 RFW ET Highlighted the influence of facial quality and racial feature gradations on fairness in face recognition models.
Skin-Tone Muthukumar et al. [36] 2018 PPB GN, ST Identified structural facial features and skin tone as key factors for dark-skinned females’ underperformance.
Cook et al. [37] 2019 Private ET, GN, AG, + Highlighted skin reflectance as a major predictor of FR disparities and emphasized acquisition methods’ role.
Lu et al. [41] 2019 IJB-B, IJB-C AG, GN, ST, + Quantified performance variations across skin tone groups, noting challenges with darker skin tones.
Krishnapriya et al. [26] 2020 MORPH ET, ST Examined FMR and FNMR across skin tones but found no direct causation between darker skin tone and higher errors.
Continued on the next page.
TABLE II: (Continued)
Reference Year Dataset Attribute Summary
Algorithmic Factors Phillips et al. [42] 2011 FRVT ET Identified the “other-race effect,” where algorithms performed better on their respective majority racial groups.
Klare et al. [27] 2012 PCSO ET, GN, AG Designed dynamic face matchers to work with models trained on exclusive cohorts.
Ricanek et al. [43] 2015 ITWCC AG Observed inefficiency in choice of algorithms towards recognizing children’s faces (due to structural changes).
Nagpal et al. [44] 2019 MORPH, RFW, CACD, + ET, AG Showed that deep learning models encode in-group biases, mirroring own-race and own-age human biases.
Albiero et al. [45] 2020 AFD, MORPH, Notre Dame GN Found gender-based biases in score distributions, with lower accuracy for women across balanced datasets.
Dooley et al. [46] 2023 CelebA, VGGFace2, RFW GN Demonstrated that network architecture and hyperparameters inherently influence fairness in FR; and bias can arise from model inductive biases.
Image Quality Lu et al. [41] 2019 IJB-B, IJB-C AG, GN, ST, + Analyzed multiple covariates; noted lighter skin tones consistently outperformed medium-dark tones.
Krishnapriya et al. [47] 2019 MORPH ET Showed improving image quality reduces performance gaps between African-American and Caucasian cohorts.
Cavazos et al. [28] 2020 GBU ET Highlighted dataset complexity and decision thresholds’ impact on racial bias and accuracy.
Albiero et al. [45] 2020 AFD, MORPH, Notre Dame GN Linked gender-based disparities to score distributions and identified confounding factors like cosmetics.
MdTF [48, 49] 2023 Private ET, GN, AG, + Found lower skin reflectance correlated with reduced accuracy and higher transaction times.
Wu et al. [50] 2023 MORPH ET, GN, + Demonstrated how brightness inconsistencies increase FMRs and diminish similarity scores.
Bhatta et al.  [51] 2024 MORPH ET, GN Studied impact of blur and resolution for 1:N identification.
Pangelinan et al. [52] 2025 MORPH ET, GN, + Balancing brightness improves similarity scores for females.
Intersectional Factors Ricanek et al. [43] 2015 ITWCC AG Discussed recognition challenges due to structural changes in children’s facial features over time.
El Khiyari et al. [53] 2016 MORPH ET, GN, AG Observed lower face verification accuracy in younger individuals, females, and Black racial groups.
Best-Rowden et al. [54] 2017 LEO_LS, PCSO_LS AG, GN Found that males generally have higher genuine scores, but their performance declines faster with age.
Vera-Rodriguez et al. [55] 2019 VGGFace2 GN Highlighted gender as a covariate, with males consistently outperforming females across demographics.
Albiero et al. [56] 2020 MORPH ET, AG Observed lack of consistent relation between recognition accuracy and age.
FRVT report [19] 2021 Private ET, GN, AG, + Noted elevated false positives for children and elderly, especially among Asian and American Indian groups.
Atzori et al. [57] 2022 VGGFace2, DiveFace ET, GN, + Explanatory framework for evaluating the influence of both protected and non-protected attributes.
Sarridis et al. [58] 2023 RFW ET, GN, AG Reported high error rates for African females over 60 years, highlighting compounded biases.
Cook et al. [49] 2023 Private ET, GN, AG, + Demonstrated that age and skin lightness significantly influence recognition scores, compounded by illumination.
Soft-Biometric Attr. Albiero et al. [59] 2022 MORPH, MFAD Face region Studies impact of facial regions on bias patterns.
Bhatta et al. [60] 2023 MORPH, MFAD Hairstyle Noted that hairstyle balance reduces gender bias.
Pangelinan et al. [61] 2024 Various Face regions Higher FMR for females despite controlling other factors.
Ozturk et al. [62] 2024 MORPH Facial hair Study of facial hairs on impostor/genuine scores.
Wu et al. [63] 2024 MORPH, BA-test Facial hair Analysis of facial hairs (occlusions) for recognition accuracy.
Mamede et al. [64] 2024 RFW (occluded) Occlusions Reported that occlusions amplify racial bias.
Kurz et al. [65] 2025 MAAD-Face Facial Attributes Gender gap vanishes with matched attributes.
III-C Algorithmic or Model Factors
While majority of sources of performance differences in FR can be traced back to data, several studies have identified how specific algorithm-related choices can amplify or introduce demographic disparities. These algorithmic elements often interact with other factors/ attributes, making causal attribution challenging. In this section, we examine works where algorithm-level factors are prominently discussed as contributors to observed bias. Phillips et al. [42] identified the “other-race effect,” where algorithms developed in Western and East Asian contexts demonstrated superior performance for their respective majority racial groups. The said effect refers to an observation that humans recognize faces of their own race more accurately than faces of other races. This disparity persisted even when datasets were balanced, hypothesizing that underlying biases may occur due to algorithmic design and training processes. However, it may be noted that most of these findings are based on systems developed over a decade ago, and their implications may vary for modern deep learning-based FR models. Klare et al.  [27] observed recognition challenges for specific demographic groups, including females, Black individuals, and younger cohorts. They reported improved performance when models were trained exclusively on specific demographic cohorts. This improvement was supported by the integration of dynamic face matchers within the recognition pipeline–a modification that reflects both algorithmic adaptation and the influence of training cohort composition. Similarly, several recent works proposed modifying features or decision thresholds based on group information (see Sec. VI). While such approaches might lead to improved accuracy for under-represented groups, it requires reliable access to demographic labels. This introduces few drawbacks: external labeling increases operational cost and may be prohibited under privacy or anti-discrimination regulations (e.g., GDPR). Alternatively, predicting demographic attributes automatically introduces another layer of uncertainty, where misclassification can directly impact system fairness and performance.
In [44], Nagpal et al. demonstrated that deep learning models encode in-group biases, mirroring human tendencies such as own-race and own-age effects. By analyzing activation maps, they showed that these biases were ingrained within the feature representations of the models. Based on these observations, they hypothesized that classification models encode faces from different racial groups in distinct ways, indicating the presence of bias. Although the study did not isolate algorithmic design as the sole cause, it highlighted how internal model behavior may reflect or reinforce demographic disparities. Dooley et al. systematically evaluated the impact of network architectures and hyper-parameters on fairness in FR [46]. Their findings revealed that architectural choices significantly influence demographic disparities. It also demonstrated that bias can be inherently encoded in a model’s inductive bias, and relying solely on standard architectures optimized for accuracy may compromise fairness.
Further investigations into gender-based disparities by Albiero et al. [45] revealed skewed impostor and genuine score distributions resulting in a lower accuracy in women. This disparity persisted across datasets, regardless of balanced training and neutral facial expressions, and hence, the authors speculated that the cause could be more intrinsic factors between the groups. Ricanek et al. [43] noted unique challenges in recognizing children’s faces due to structural changes with age, finding that algorithms effective on adult faces performed poorly for younger subjects. Although this issue pertains to longitudinal variations in facial morphology, it underscores a broader implication: the same algorithms may not be equally effective across seemingly similar tasks (FR). Attributes such as facial structure should therefore be considered as critical factors in guiding the design and selection of appropriate recognition algorithms.
Identifying algorithmic factors that contribute to differential recognition accuracy across demographic groups remains a complex task. This process requires isolating the influence of the training algorithm–an integral component of the recognition pipeline. To ensure valid attribution, experimental setups need to be carefully controlled that differ in nothing but algorithms. Initially, performance disparities were primarily attributed to handful factors such as imbalances in training datasets or variations in acquisition conditions. However, as the number of potential influencing factors grows, the experimental studies to rigorously investigate and isolate the algorithmic sources of performance differences require a thorough and systematic approach.
III-D Image Quality
The quality of input images and associated covariates significantly influence the manifestation of demographic differentials in FR systems. Numerous studies have emphasized how disparities in image quality across different demographic groups can lead to variations in system performance. For instance, Cavazos et al. [28] analyzed both data-driven– focusing on aspects of image quality– and scenario-based factors, revealing that dataset complexity and decision thresholds have a notable impact on recognition accuracy and racial bias. Their experiments across multiple algorithms further demonstrated that East-Asian faces required higher decision thresholds compared to Caucasian faces to achieve equivalent error rates. The authors also acknowledged that image quality may not act in isolation and that multiple underlying causes may jointly contribute to observed disparities.
The study conducted by MdTF highlighted skin reflectance as a critical factor influencing both the accuracy and efficiency of FR systems [48, 49]. Analyzing 158 FR systems, they found that lower skin reflectance, typically associated with darker skin-tones, correlated with reduced accuracy and higher transaction times. These effects varied across systems, underscoring the role of image acquisition quality as a stronger predictor of performance as mentioned earlier. Similarly, Wu et al. [50] explored the effects of brightness and illumination, demonstrating that under-exposed or over-exposed images result in higher FMRs, while significant brightness differences between image pairs diminish similarity scores. They recommended controlled image acquisition processes to achieve consistent brightness across demographic groups, thereby reducing accuracy disparities. In [52], authors analyzed the effect of facial brightness on recognition performance disparities between Caucasian and African American females. Their experiments showed that by balancing brightness–using either median pixel values or full pixel distributions–the similarity scores improve for both groups.
Krishnapriya et al. [47] further examined how variations in image quality contributed to performance gaps between African-American and Caucasian cohorts. Enhancing image quality notably reduced these disparities, particularly by minimizing low-similarity errors within the genuine distribution. Following ICAO compliance guidelines, they evaluated biometric sample quality to support these findings. Albiero et al. [45] investigated gender-based disparities in FR systems, linking these to differences in genuine and imposter score distributions. They also identified confounding factors such as cosmetics and image pose. Despite using neutral and balanced datasets, their study revealed that such measures alone were insufficient to fully eliminate observed disparities. Bhatta et al.  [51] explored 1-to-many identification performance under degraded image conditions resembling surveillance-quality probes. The findings showed that blur and low resolution significantly increased false matches, with demographic disparities being more pronounced between genders than racial groups, underscoring the influence of image quality on recognition fairness. The investigation of demographic fairness in the context of 1-to-many (1:$N$) face identification has been comparatively under-explored. Another recent study in area [66] observed that the accuracy variation patterns in 1:$N$ identification are similar to those reported in 1:1 verification scenarios. Furthermore, their analysis highlighted that degradations in image quality–particularly due to blur or reduced resolution–tend to increase the false positive identification rate (FPIR).
Work by Lu et al. provided a detailed analysis of the influence of covariates on FR performance, incorporating variables such as skin-tone, age, gender, pose, facial hair, and occlusion across three datasets and five FR systems [41]. Their findings highlighted that skin tone significantly affects verification accuracy, with lighter skin tones consistently outperforming medium-dark tones. However, they also emphasized the challenges posed by ambiguities in skin-tone classification, advocating for more precise methodologies for performance assessments. In alignment with earlier studies, Lu et al. observed that male subjects generally achieved better recognition accuracy than female subjects. They attributed this disparity to factors such as occlusion caused by longer hair and alterations in facial appearance due to makeup. These observations corroborate prior findings indicating that facial makeup can negatively impact recognition accuracy [67, 68]. Collectively, these studies underscore that demographic fairness in FR systems is intrinsically tied to image quality and related covariates, necessitating focused efforts to address these issues systematically.
III-E Combined or Intersectional Factors
In the preceding sections, we examined the individual factors contributing to demographic differentials in FR. This section shifts focus to studies that investigate the combined effects of multiple demographic attributes, such as age, race, and gender. Vera-Rodriguez et al. [55] emphasized the significance of gender as a covariate in FR, observing that males consistently outperform females across various demographic groups. These findings highlight the necessity of addressing combined demographic factors to achieve equitable outcomes.
Age-related disparities in FR performance have been linked to structural transformations in facial features over time, particularly among children. Ricanek et al. [43] observed increased complexity of child aging compared to adults, attributing recognition challenges to changes in facial bone structure and the proportions of facial components. Additionally, Best-Rowden and Jain [54] reported nuanced patterns in age-related recognition performance, noting that while males generally exhibit higher genuine scores, their performance declines more rapidly with age compared to females. These observations underline the intricate interplay of demographic attributes in shaping performance of FR systems. The intersection of age, race, and gender significantly amplifies accuracy differentials in FR systems. In [58], Sarridis et al. identified a disproportionately high number of incorrect verifications (both false positives and false negative)for African females over 60 years compared to Caucasians, illustrating the compounded effects of intersecting demographic factors. Similarly, El Khiyari et al. [53] demonstrated that face verification accuracy is notably lower for younger individuals (aged 18–30), females, and certain racial groups such as Black individuals, highlighting the challenges posed by such intersections of demographic factors. Liu et al.  [69] observed that multiple factors such as ethnicity, pose, occlusion, and blur jointly impact FR accuracy.
Among the few studies addressing explanatory frameworks, [57] introduced a methodology aimed at evaluating the influence of both protected and non-protected attributes on FR performance. Here, protected attributes refer to human characteristics legally safeguarded from discrimination (e.g., race or gender), whereas non-protected attributes include modifiable traits such as makeup or hairstyle. Their analysis revealed that fairness outcomes exhibit considerable variability across multi-attribute groupings. Notably, patterns observed in single-attribute evaluations frequently diminish or invert when extended to analyses involving multiple attributes.
Algorithmic evaluations further reinforce these findings. The FRVT report [19] observed elevated false positives among children and the elderly, particularly within Asian and American Indian groups. These disparities were intensified in low-quality imaging conditions, with younger and older demographics experiencing higher error rates. Cook et al. [49] extended this understanding by showing that self-reported demographic factors like age and measured skin lightness significantly impact recognition scores, often compounded by environmental factors such as illumination. However, more recent analyses, such as Albiero et al.  [56], suggest that recognition accuracy may not consistently improve or decline with age. Their findings indicate that face matchers do not exhibit a simple age-performance trend, challenging assumptions from earlier works.
The inconsistent findings indicate that age-related performance differences in FR remains insufficiently understood and lacks a clear consensus. Overall these studies highlight the complexity of mitigating intersecting biases within FR systems, underscoring the importance of incorporating age-specific and multi-factor considerations in both algorithm development and evaluation protocols.
III-F Soft-Biometric Attributes or Non-Demographic Factors
While prior sections have explored to demographic factors influencing disparities in FR performance, recent research has highlighted the critical role of non-demographic factors– such as facial hair, hairstyle, and makeup– in shaping recognition outcomes. These factors, though not inherently demographic, are deeply intertwined with social and cultural norms that vary across gender and ethnicity. In the context of FR, these soft attributes function as partial occlusions or lead to shifts in the underlying data distribution, thereby introducing additional challenges. It is well-documented that occlusions can significantly degrade recognition accuracy in biometric and generic vision applications [70, 71, 72, 73]. However, when these occlusions vary systematically across groups, they contribute to unequal recognition outcomes and further exacerbate existing disparities, mimicking demographic bias.
Recent studies have demonstrated that many of the observed demographic fairness in FR may in fact be driven by these correlated non-demographic traits. Through extensive experiments, Kurz et al. showed that the gender gap in recognition accuracy vanishes when males and females share the same facial attributes, indicating that social appearance conventions, rather than gender itself, are a primary factor [65]. Pangelinan et al. similarly found that controlling for facial pixel information across genders improved non-matching errors for females, pointing to the role of hairstyles and other aesthetic choices in accuracy differentials [61]. Multiple works have rigorously examined the impact of facial hair, showing that variations in facial hair significantly affect both impostor and genuine score distributions, and that these effects persist even with balanced training datasets[62, 63]. In females, factors such as hairstyle, facial morphology, and overall facial dimensions result in a reduced number of effective face pixels, which has been associated with elevated false match rates (FMR) compared to males [74]. Bhatta et al.  [60] demonstrated that balancing for hairstyle-induced occlusions in test data can nearly eliminate the gender gap. Additionally, Albiero et al. illustrated that recognition performance can vary across different facial regions, further suggesting that visual composition and occlusions, rather than demographic identity alone, play a pivotal role toward recognition [59]. Complementing these findings, Mamede et al. [64] proposed the Face Occlusion Impact Ratio (FOIR)- a novel metric to measure the influence of occlusions on recognition performance across demographic groups. Their results show that occlusions impact demographic groups unequally, with African individuals disproportionately affected- thereby amplifying existing biases.
These collective insights underscore the significant influence of non-demographic but demographically correlated appearance factors in shaping recognition performance. While these factors are not inherently biological, but rather linked to the demographic groups through social and cultural appearance norms and styles. Understanding and accounting for these nuanced sources of variation is essential to mitigate differential accuracy and build fair FR systems.
IV Datasets for the Study of Demographic Accuracy Differences
In evaluating and mitigating demographic fairness in FR systems, the selection of suitable datasets plays an important role. Although numerous FR datasets exist, those specifically intended for differential performance and fairness-related tasks must include demographic labels associated with each subject or identity. The datasets designed for tasks such as race, gender, ethnicity, or age estimation are particularly useful when they include demographic labels, as such factors are critical for assessing fairness. For certain tasks (related to estimation or classification of attributes), having a single image per subject may suffice, as training and testing can be conducted separately for each image. However, FR models, especially state-of-the-art systems, benefit from having multiple images (variations) of each identity to train more robust feature extractors. The testing phase or deployment has two primary modes: one-to-one (1:1) verification and one-to-many (1:$N$) identification. In 1:1 matching, a probe image is compared against one or more enrolled images of a claimed identity. For identification (1:$N$ matching), the probe is compared against a gallery containing reference images from multiple identities to determine the closest match. Occasionally, the gallery may be composed of multiple images per identity.
TABLE III: Commonly used datasets for tasks related to demographic accuracy differentials in face recognition.
Dataset Name Year # Images / Subjects Demographic Labels Typical Purpose
MORPH-II [75] 2008/2016 55,000 / 13,000 (Male, Female), (Black, White, Asian, Hispanic) Train/ Test
MEDS-II [76] 2011 1,300+ / 518 (Male, Female), (Asian, Native American, Black, White) Test
Adience [77] 2014 26,580 / 2,284 (Male, Female) Test
AFD (Curated) [78][79] 2018 91,000+ / 1,878 (Male, Female), Asian Train
VGGFace2 [80] 2018 3.31M / 9,131 (Asian, Black, Indian, White) Train/ Test
DemogPairs [81] 2019 10,800 / 600 (Male, Female), (Asian, Black, White) Test
RFW [34] 2019 24,000 pairs / - (Asian, African, Caucasian, Indian) Test
KANFace [82] 2020 40 K (+ 44 K videos) / 1,045 (Male, Female) Test
BUPT-BalancedFace [83] 2020 1.3M / 28,000 (Asian, African, Caucasian, Indian) (7K per race) Train
BUPT-GlobalFace [83] 2020 2M+ / 38,000 (Asian, African, Caucasian, Indian) Train/ Test
DiveFace [84] 2020 120,000 / 24,000 (Male, Female), (Asian, African, European) Train/ Test
BFW [85] [86] 2020 20,000 / 800 (Male, Female), (Asian, Black, Indian, White) Test
MORPH-III [74] 2021 127,000+ / 26,401 (Male, Female), (Black, White) Train/ Test
CASIA-Face-Africa [87] 2021 38,500+ / 1183 (Male, Female), African Train/ Test
These requirements significantly reduce the availability of datasets suitable for assessing demographic fairness, as most FR datasets do not provide adequate demographic labels or have highly skewed distributions of subjects across different demographic groups. It is important to note that several prominent bias evaluation efforts, such as those led by NIST and the MdTF, rely on proprietary or restricted-access datasets. While these organizations have conducted large-scale, operationally relevant evaluations, their limited availability to the wider research community poses challenges for reproducibility, benchmarking, and comparative analysis. In this section, we outline recent and commonly used publicly available datasets that are relevant for assessing and addressing demographic fairness in FR systems. The datasets presented below are arranged in chronological order of their release.
  • MORPH [75]: The MORPH dataset is one of the largest facial image datasets available in several variants and versions. MORPH-II is the most commonly referred academic version, comprising more than 55,000 images from more than 13,000 subjects. Despite its usefulness, it should be noted that the dataset is highly skewed in terms of gender and ethnicity, with a significant over-representation of male subjects (more than 46,000 images) and a limited number of female subjects (approximately 8,500 images).
    The MORPH-III dataset has seen increased adoption in recent research investigating demographic variations in FR [74]. For improved data quality, the dataset was curated to remove duplicates, twins, and mislabeled entries, resulting in a refined set comprising 127,000+ images of nearly 26,000 subjects. This curated subset provides a cleaner and demographically structured resource for evaluating bias. For both versions, the datasets consist of mugshot-style facial images acquired under controlled conditions, featuring neutral expressions, approximately frontal poses, and consistent illumination. However, it has been noted to have demographic imbalances, particularly under-representation of certain groups (e.g., Asian, Indian), which may affect fairness studies.
  • Multiple Encounter Dataset (MEDS-II) [76]: The MEDS-II dataset is an extension of the MEDS-I dataset and was created to assist with the NIST Multiple Biometric Evaluation. The dataset includes over 1,300 images of 518 subjects, with many subjects having only a single image, limiting its usefulness for verification tasks. The MEDS-II dataset is dominated by male subjects of White and Black ethnicities. Despite its limitations in demographic diversity, it remains a useful resource for testing FR systems in real-world scenarios, especially where multiple encounters of a subject are available. Additionally, the dataset is considerably smaller compared to current benchmarks.
  • Adience [77]: The Adience dataset consists of 26,580 images from 2,284 individuals captured in unconstrained (in the wild) conditions using typical consumer devices. The dataset is annotated with identity, gender, and one of eight age group labels. While primarily designed for age and gender classification tasks, the dataset’s natural variation in pose, illumination, and demographics makes it valuable for FR research. Sourced from nearly 200 publicly available Flickr albums, this dataset has a higher representation of female subjects.
  • AFD (Asian Faces Dataset - Curated) [78]: The Asian Faces Dataset (AFD) was developed using images scraped from the web, with a focus on frontal face images [79]. The curated version, provided by Zhang et al.  [78] includes over 42,000 images of 911 males and 49,000 images of 967 females. A gender classifier was used to filter out mislabelled images, and duplicate or near-duplicate images were removed. This curated dataset is useful for studying gender and ethnic bias in FR systems, specifically for models focused on the Asian demographic.
  • VGGFace2 [80]: The VGGFace2 dataset is a large-scale FR dataset containing over 3.31 million images of 9,131 subjects. This dataset was annotated for gender (Male, Female) and ethnicity (Asian, Black, Indian, White) labels by Idiap Research Institute [^1] Annotations for VGGFace2 Dataset making it useful for fairness related tasks. This dataset offers significant advantages due to a large number of identities with variations in age, pose, and lighting; VGGFace2 improves over the original with more diversity too. The dataset, however, is significantly imbalanced, with a disproportionate representation of White and male subjects. Although this dataset offers high intra-class variation with over 3 million images, it contains only slightly more than 9,000 unique identities, which may limit its utility for evaluating false positive rates in large-scale evaluations. Note that, the original download source has been removed by the creators as of 2024.
  • DemogPairs [81]: The DemogPairs dataset is a validation set containing 10.8K images, divided into six demographic folds: Asian females, Asian males, Black females, Black males, White females, and White males. The dataset was specifically designed to evaluate the demographic disparities in FR models, offering 58.3 million evaluation pairs, including cross-demographic, cross-gender, and cross-ethnicity pairs. The DemogPairs dataset was constructed with rigorous demographic annotation and is a useful resource for testing the generalization of FR systems across diverse demographic groups.
  • Racial Faces in-the-Wild (RFW) [34]: The RFW dataset is a benchmarking dataset designed to study racial fairness in FR systems. It consists of four subsets (African, Asian, Caucasian, and Indian), each containing about 3,000 individuals and 6,000 image pairs (these pairs have been defined by its creators). The dataset is specifically used for face verification tasks and includes balanced pairs of genuine (mated) and imposter (non-mated) images. The RFW dataset has been widely adopted by the research community to evaluate and compare the performance of FR algorithms across different racial groups. However, given the overlap between the RFW dataset and the widely used training dataset, MS-Celeb-1M, it is important to consider that the evaluation set may have been seen by the model during training. The creators of RFW have also provided a list of non-overlapping images to address this issue.
  • KANFace [82]: The KANFace dataset comprises around 40,000 still images and 44,000 video clips (totaling roughly 14.5 million frames) of 1,045 individuals, collected in unconstrained settings. On average, each subject is represented by 39 images and over 13,000 video frames. The dataset includes manual annotations for identity, age, gender, and kinship, with demographic metadata sourced from Wikipedia and IMDb. The dataset consists of 586 male and 459 female subjects.
  • BUPT-BalancedFace [83]: The BUPT-BalancedFace dataset was constructed to address demographic fairness by ensuring race balance across the dataset. It contains approximately 1.3 million images from 28,000 celebrities, with a balanced distribution of 7,000 identities per race. The dataset was selected from MS-Celeb-1M dataset [88] through the FreeBase and Face++ APIs, although it has been noted that the labels may contain noise. Due to its size and balanced nature, the BUPT-BalancedFace dataset has become a popular resource for training and fine-tuning FR models while mitigating race-related biases. Two key points should be considered: first, the demographic labels, which are based on nationality or API, contain some noise. Second, the dataset overlaps with MS-Celeb-1M dataset, which restricts certain experiments when using models trained exclusively on MS-Celeb-1M dataset.
  • BUPT-GlobalFace [83]: In addition to the BUPT-BalancedFace dataset, the BUPT-GlobalFace dataset is also widely used in fairness tasks. Unlike its balanced counterpart, BUPT-GlobalFace is racially imbalanced, containing over 2 million images from approximately 38,000 celebrities, with an over-representation of certain racial groups (e.g., Caucasian and Asian). Although it reflects real-world distribution more closely, the imbalance makes it less suitable for bias mitigation training and more appropriate for evaluating bias under unconstrained, skewed demographic settings.
    Figure 3: Distribution of images of commonly used FR datasets considering (a) race and (b) gender as demographic factors. The details of distribution have been used from original sources (wherever available) or from other works contributing to this information; while the naming convention has been altered for unified representation aligning to the convention used by most datasets.
  • DiveFace [84]: The DiveFace is a dataset generated from the Megaface dataset (now decommissioned) [89], containing over 120,000 images from 24,000 identities. The dataset includes two gender and three ethnicity classes, allowing for detailed demographic analysis. Annotations were made using a semi-automatic process, followed by manual inspection. This dataset is useful for studying the impact of gender and ethnicity in FR tasks, although it may exhibit bias in some groups. It should also be noted that in DiveFace dataset, the subjects of Indian and African ethnicities have been grouped together– which can make it difficult to use it in conjunction with other datasets that typically do not follow such grouping.
  • BFW (Balanced Faces in the Wild) [85] [86]: The BFW dataset was designed to provide a more balanced evaluation of FR systems by creating subgroups that are evenly split across gender and ethnicity. The dataset is compiled from VGGFace2 [80] and offers a refined approach to subgroup analysis with less overlap between training and testing data. The corresponding demographic labels were generated using ethnicity [90] and gender [91] classifiers, followed by manual validation. Additionally, the BFW dataset is also balanced with respect to the number of images, subjects, and the (ratio of) images per subject; making it particularly useful for evaluating the demographic fairness of FR models, offering a more balanced alternative to other datasets. However, this dataset shares the same two limitations as some earlier ones: a smaller size and overlap with VGGFace2.
  • CASIA-Face-Africa [87]: The CASIA-Face-Africa dataset is the first large-scale face dataset of African subjects– comprising 38,546 images from 1,183 individuals, captured under varying illumination conditions using multi-spectral cameras. It includes detailed demographic attributes and facial expressions along with manually annotated with facial key points. The dataset exhibits a well-distributed age representation, with a significant portion belonging to the subjects up to 40 years, aligning with the majority workforce demographics. Additionally, it maintains an almost balanced gender ratio (48% male, 52% female), making it useful for gender-based analysis as well. In terms of ethnic variations, the dataset includes multiple African ethnic groups, with a notable dominance of the Hausa ethnic group.
The datasets are summarized in Table III, and Fig. 3 illustrates an overview of the demographic distribution, with race and gender as demographic variables (where details of both variables were available), for the datasets reviewed in this section.
V Assessment of Demographic Fairness in FR
The bias in FR system leads to variations in score distributions and their overlaps across different demographic groups. Such disparities inherently result in different FMR and FNMR values for each group when a single (global) threshold is used. Fig. 1 illustrates score distributions of different groups may differ, leading to unequal error rates. In this section we briefly review performance measures designed for demographic-aware assessments of fairness in FR systems. Given the overlap and distinct characteristics of scores and decisions, establishing a well-defined evaluation framework is crucial for accurate analysis. Howard et al.  [10] introduced the concepts of differential performance and differential outcomes, providing two key terms that aid in achieving a precise understanding and categorization of assessment metrics.
  • Differential Performance: Variations in genuine (mated) or imposter (non-mated) distributions across demographic groups, independent of thresholds.
  • Differential Outcome: Differences in FMRs or FNMRs between groups, based on decision thresholds.
Quantifying demographic bias or fairness, in terms of both demographic performance and outcomes, is critical for developing fair and reliable FR systems. By analyzing the overlap of distributions and error disparities, one can identify specific areas requiring intervention to ensure equity across diverse user groups.
Since the 2019 Face Recognition Vendor Test (FRVT) [^2] Since 2023, the FRVT initiative has been restructured into the Face Recognition Technology Evaluation (FRTE) and Face Analysis Technology Evaluation (FATE) programs. report, NIST has included demographic effects in FR algorithms [19]. This assessment involved comparing non-mated pairs within the same demographic group, setting thresholds for algorithms to achieve an FMR of 0.0001 for white males (since this demographic typically associated with the lowest FMR). Constraining non-mated pairs to share the same demographic attributes (e.g., race or gender) is commonly referred to as yoking [92]. The report [19] offered a comprehensive analysis of recognition processes and identified areas where demographic effects might occur. To quantify demographic disparity, NIST initially employed Inequity Ratios (IR), calculating the ratio of maximum to minimum of FMR and FNMR, at score threshold $\tau$, across demographic groups as shown in Eq. 1.
$\text{IR}(\tau)=\left(\frac{\max_{d_{i}}\text{FMR}_{d_{i}}(\tau)}{\min_{d_{j}}\text{FMR}_{d_{j}}(\tau)}\right)^{\alpha}\times\left(\frac{\max_{d_{i}}\text{FNMR}_{d_{i}}(\tau)}{\min_{d_{j}}\text{FNMR}_{d_{j}}(\tau)}\right)^{\beta},$ (1)
–where, the subscripts $d_{i},d_{j}$ refer to the demographic group; and $\alpha,\beta$ are weighing coefficients. The error rates (FMR and FNMR) have been yoked, meaning they were calculated using pairs exclusively from the same demographic group.
However, considering potentially large range of the error rates, these ratios can become numerically unstable, especially in extreme cases. To alleviate this shortcoming, NIST has considered few modified versions of Inequity Ratios such as adjusting the score threshold ($\tau$), incorporating fixed constants in the denominator, or expressing worst-case error rates relative to arithmetic or geometric means [93]. Using the geometric mean is particularly advantageous due to its extended range over FMR/FNMR values, and this metric, as described in Eq 2 has often been used by NIST.
$\text{A}(\tau)=\frac{\max_{d_{i}}\text{FMR}_{d_{i}}(\tau)}{\text{FMR}^{*}(\tau)}$
$\text{B}(\tau)=\frac{\max_{d_{i}}\text{FNMR}_{d_{i}}(\tau)}{\text{FNMR}^{*}(\tau)},$ (2)
–where the error rates marked with represent the geometric mean computed across all demographic groups. In [94], the authors computed the ratio of each group’s error rate (FMR or FNMR) to the geometric mean across all demographic groups. The overall bias metric was then defined as the sum of the logarithms ($\log 10$) of these ratios for all groups. Another possible approach involves referencing a standard FMR or FNMR in the denominator, which inherently resolves stability issues while providing more robust evaluations of demographic fairness.
It is important to note that the IR metric defined in Eq.1 combines two components: one derived solely from the group-wise FMR values and the other from FNMR values. While the metric multiplies these components after applying user-defined weights ($\alpha$, $\beta$), it is also possible to use the individual terms independently (either or both). The constituent ratios (e.g., the max to min FMR or FNMR across demographic groups, in this case) provide interpretable measures of disparity. The choice of weights typically depends on the specific requirements of the end application. Several prior works have evaluated fairness using single-component metrics focusing on either FMR or FNMR separately[19, 95]. Most metrics, that assess difference in FR performance, discussed below also derive from these two fundamental error types. For brevity, we focus on combined formulations, but note that using one-sided metrics (which is also equivalent to setting one weight to minimum permissible value) is also a valid and reported practice when appropriately motivated.
The Fairness Discrepancy Rate (FDR) is one of the initial efforts of quantifying fairness in FR [96]. For an FR system using a single decision threshold, the FDR combines FMR and FNMR using weighted sums, and evaluates fairness through a unified measure that captures the trade-offs between both error rates. As described by Eq. 3, the FDR requires two hyper-parameters: one for the score threshold ($\tau$), and another for defining relative importance ($\alpha$) of FMR over FNMR. Thus, it offers flexibility of assessing the fair nature of the model at pre-defined score threshold and application-dependent weighing of false matches to false non-matches.
$A(\tau)=\max\left|\text{FMR}_{d_{i}}(\tau)-\text{FMR}_{d_{j}}(\tau)\right|,$
$B(\tau)=\max\left|\text{FNMR}_{d_{i}}(\tau)-\text{FNMR}_{d_{j}}(\tau)\right|,$
$\text{FDR}(\tau)=1-\left(\alpha A(\tau)+(1-\alpha)B(\tau)\right).$ (3)
In a similar vein, to define the balance between false matches and false non-matches, NIST proposed an approach to calculate the Inequity Measure by raising the terms representing demographic disparities to specific exponents (serving as weights) and then multiplying them.
TABLE IV: Summary of fairness assessment metrics employed in demographic analysis of face recognition systems.
Assessment Metric Ref Publication Year Description
Inequity Rate (IR) NIST / Grother et al.  [20] 2021 This metric involves the analysis of the minimum and maximum ratios of False Match Rate (FMR) and False Non-Match Rate (FNMR), with the application of weighted exponents.
Fairness Discrepancy Rate (FDR) Pereira & Marcel [96] 2021 A weighted combination of the maximum differential values of FMR and FNMR.
Statistical Approaches Schuckers et al.  [97] 2022 Utilizes a bootstrap-based hypothesis testing approach to assess bias.
Separation/Compactness Metrics Kotwal & Marcel [98] 2022 Investigates the distributions of genuine and impostor scores, focusing on shape and compactness characteristics.
GARBE (Gini Coefficient Based Metric) MdTF / Howard et al.  [99] 2022 Measures statistical dispersion in FMR and FNMR, integrating these through linear weighting.
Mean Absolute Percentage Error (MAPE) Villalobos et al.  [100] 2022 Calculates the relative deviation of FMR from a pre-established benchmark.
Sum of Group Error Differences Elobaid et al.  [101] 2024 Examines relative deviations in group-level FMR and FNMR compared to global scores, providing insight into disparities.
Comprehensive Equity Index (CEI) Solano et al.  [102] 2024 A weighted combination that assesses disparities in both tail and central distributions of performance metrics.
ROC Uncertainty Conti & Clémençon [94] 2024 Utilizes normalized uncertainty of the ROC curves to assess fairness.
Standard Deviation - - The standard deviation of FMR, FNMR, and True Match Rate (TMR) is assessed to gauge variability.
d-prime ($d^{\prime}$) - - Quantifies the separation between score distributions of different demographic groups; higher values indicate greater separability.
Skewed Error Ratio (SER) - - A ratio that compares the worst-case error rates across different demographic groups.
Mean Absolute Deviation (MAD) - - Evaluates the MAD in specific error rate computed at a fixed threshold.
Trade-Off (TO) - - Evaluates the trade-off between performance metrics, particularly the difference in average accuracy and standard deviation.
Schuckers et al. highlighted the importance of accounting for statistical variation when evaluating fairness in FR systems [97]. They noted that the differences among demographic groups can arise either from actual performance disparities or by chance due to sampling variability, leading to potential Type-I errors. To address this, they proposed two statistical methodologies: a bootstrap-based hypothesis test and a simpler test methodology tailored for non-statistical audiences. Their study also conducted simulations to explore the relationship between margin of error and factors such as the number of subjects, attempts, correlation between attempts, underlying FNMRs, and the number of demographic groups. In [94], the authors propose a statistical framework for assessing fairness using Receiver Operating Characteristics (ROC) curves and associated uncertainty measures derived from U-statistics. They argue that the ROC curve is not only central to evaluating system performance but also serves as a key tool for analyzing fairness in similarity scoring. They established asymptotic guarantees for empirical ROC curves and fairness-relevant metrics, emphasizing the importance of accurately quantifying uncertainty. Notably, they highlight that conventional bootstrap methods may misrepresent uncertainty due to the U-statistical nature of FMR and FNMR, and recommend a recentering technique to ensure valid inference. In [99], the researchers at MdTF proposed new metric for assessment of bias based on demographic outcomes. Their metric, GARBE (Gini Aggregation Rate for Biometric Equitability), is inspired by the Gini coefficient– which has a long history of use as a dispersion measure in socio-economic context. The GARBE evaluates statistical dispersion in error rates and emphasizes equitable treatment across demographic groups. Similar to the FDR and IR, this metric combines weighted contributions of FMR and FNMR to produce a single fairness score for a given fixed score threshold. The formula for GABRE and its constituent Gini coefficient are provided in Eq. 4.
GARBE $=\alpha\,G_{\text{FMR}}(\tau)+(1-\alpha)\,G_{\text{FNMR}}(\tau),\quad\text{where},$
$G_{x}$ $=\left(\frac{n}{n-1}\right)\left(\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}|x_{i}-x_{j}|}{2n^{2}\bar{x}}\right).$ (4)
Metrics such as GARBE and several others mentioned earlier provide a single scalar value to quantify fairness. This reporting has many advantages as it facilitates simplified, easy comparisons and ranking of systems. However, reducing multiple fairness indicators into a single combined measure as well introduces certain limitations. In particular, the choice of weighting coefficients is often subjective or application-dependent, and may not reflect the actual impact of each component. Moreover, combining ratios (e.g., from FMR and FNMR) that operate on different scales of magnitude can inadvertently skew the metric toward one aspect of fairness, thereby distorting the overall assessment. If the weights are not thoughtfully chosen and justified, such aggregation may obscure significant disparities between demographic groups.
Villalobos et al. proposed the Mean Absolute Percentage Error (MAPE) as a metric to quantify differences in error rates across demographic groups [100]. MAPE measures the relative deviation of FMRs from a policy-defined FMR, ensuring that low error rates for one group do not mask higher error rates for another. High deviations in error rates, particularly towards lower values of FMR, can negatively impact the system by increasing FNMR. A MAPE score of zero indicates that all demographic groups achieve the desired FMR, making it an effective metric for fairness evaluation. If $\text{FMR}_{p}$ indicates policy-defined FMR, then MAPE for a system using $N$ groups can be computed as:
$\text{MAPE}=\frac{100}{N}\,\sum_{d=1}^{N}\left|\frac{\text{FMR}_{d}-\text{FMR}_{p}}{\text{FMR}_{p}}\right|.$ (5)
The Sum of Group Error Differences (SEDG) was introduced as the fairness assessment metric address disparities in biometric verification systems in [101]. The SEDG calculates relative deviations in FMR and FNMR across demographic groups from the FMR/ FNMR of global scores. They consider the Equal Error Rate (EER) threshold as a reference to compute the error rates. In other words, it adapts a relative difference formula to quantify demographic fairness by comparing individual group performances to a global standard. Authors argue that by incorporating both within-demographic (WDI) and cross-demographic (CDI) interactions, SEDG is able to provide better understanding of the magnitude and type of bias making it a versatile measure. For a setup comprising $\mathcal{D}$ set of demographic groups, the SEDG can be obtained as:
$\delta X_{d}=\left|1-\frac{X_{d}(\tau)}{X_{\text{global}}(\tau)}\right|,\cdots\quad X\in(\text{FMR},\text{FNMR}),$
$\text{SED}_{\rm G}=\{\delta\text{FMR}_{d}+\delta\text{FNMR}_{d}\,\,|\,d\in\mathcal{D}\}.$ (6)
Relatively fewer attempts have been made to assess the demographic fairness at score-level (i.e., based on differential performance). Among the various metrics used to quantify fairness in biometric systems, d-prime ($d^{\prime}$) has seen increasing adoption [51, 60, 50]. Originally from signal detection theory, it measures the distance between score distributions of mated and non-mated pairs, and can be adapted to assess separability between demographic group distributions. It is threshold-independent, making it robust in cases where system decision thresholds are not aligned across deployments. If $\mu,\sigma$ represent the mean and standard deviations of the scores, the $d^{\prime}$ is obtained as:
$d^{\prime}=\frac{|\mu_{m}-\mu_{nm}|}{\sqrt{\frac{1}{2}(\sigma_{m}^{2}+\sigma_{nm}^{2})}},$ (7)
—where the subscripts $m$ and $nm$ refer to the distributions of mated and non-mated pairs, respectively. The Kolmogorov–Smirnov (KS) bias metric has been adapted for assessing demographic bias in FR by measuring disparities between the score distributions of different demographic groups [103].
In [98], Kotwal and Marcel introduced three fairness evaluation measures that emphasize the separation, compactness, and distribution of genuine and impostor scores. Unlike conventional approaches that depend on system accuracy, these measures focus on assessing differential performance without requiring external parameters such as score thresholds. By examining how well the match is, rather than merely determining a match, this approach provides a more nuanced evaluation of demographic fairness. The Separation Fairness Index (SFI) measures the consistency in separation between the mean genuine and impostor scores across groups–higher consistency implies better fairness. The Compactness Fairness Index (CFI) assesses the similarity in intra-group score variability by comparing the standard deviations of genuine and impostor scores across demographics. The Distribution Fairness Index (DFI) evaluates how closely the entire score distribution of each group matches a reference distribution (e.g., the population average), using a statistical divergence measure. Their normal variants are summarized in Eq. 8.
$S(d)=|\mu_{m}(d)-\mu_{nm}(d)|;\text{and }C(d)=\sigma_{m}(d)+\sigma_{nm}(d);$
$\text{SFI}=1-\frac{2}{N}\sum_{d=1}^{N}\left|S(d)-\bar{S}\right|;\quad\text{CFI}=1-\frac{2}{N}\sum_{d=1}^{N}|C(d)-\bar{C}|$
$\text{DFI}=1-\frac{1}{N\,\log_{2}N}\sum_{d=1}^{N}\mathcal{D}_{\rm KL}\left(\mathbb{P}(d)\,||\,\mathbb{P}_{\text{ref}}\right)$ (8)
Here, $N$ denotes the number of demographic groups; $\mu_{m}(d)$ and $\mu_{nm}(d)$ are the mean mated and non-mated scores for group $d$; $\sigma_{m}(d)$ and $\sigma_{nm}(d)$ are their standard deviations; $\mathbb{P}(d)$ is the score distribution for group $d$; $\mathbb{P}_{\text{ref}}$ is the reference distribution (often the average across groups); and $\mathcal{D}_{\rm KL}$ represents the KL divergence. Building on the work from [98], Solano et al. developed the Comprehensive Equity Index (CEI)– combining error rate differences and recognition score distribution disparities [102]. The CEI enhances bias quantification methods by considering both the distribution tails and overall shapes of score distributions, enabling the detection of subtle biases across demographic groups. They also conducted experiments on high performing FR systems (as per NIST evaluations) using challenging, real-world datasets. Their experiments showed that CEI was able to effectively capture the demographic bias on several challenging datasets with several covariates. In compact manner, the CEI can be described as: $\text{CEI}=w_{1}\cdot\text{Tail Disparity}+w_{2}\cdot\text{Central Disparity}$. The disparity calculation is similar to that of DFI expressed in Eq. 8.
Several studies have evaluated bias and fairness in FR systems using the standard deviation of performance metrics calculated across demographic groups. These metrics include FMR, FNMR, and True Match Rate (TMR)– where higher value in standard deviation corresponds to greater demographic disparities [104, 105, 41, 86, 83, 100]. The mean absolute deviation (MAD) of a metric across demographic groups is another approach used to quantify bias. For example, in [106], Terhörst et al. evaluated bias by computing the MAD of the TMR at a fixed FMR threshold. Another significant metric is the Skewed Error Ratio (SER), which specifically focuses on worst-case error ratios, providing insights into the performance imbalance across groups [100, 83, 107]. Recent competitions [108, 109, 110] exploring the use of synthetic data for FR and bias mitigation have adopted a trade-off performance metric: the mean accuracy adjusted by the standard deviation. This metric aims to ensure that efforts to mitigate bias do not come at the expense of recognition performance. This metric emphasizes the development of FR models that achieve both high recognition performance and fairness across constituent demographic groups.
Table IV provides a brief summary of the fairness assessment metrics discussed in this section. While several metrics listed here capture similar goals: quantifying inter-group performance disparities, they differ in sensitivity, interpretability, and need for threshold selection. Some, that employ differential FMR/FNMR, are easier to relate to operational performance but may not generalize across systems with differing threshold settings. Others, like d-prime and fairness index measures, focus on distributional shifts and are less sensitive to specific thresholds. An important yet often under-addressed aspect in fairness evaluation is the choice of threshold setting in threshold-based metrics. Most studies adopt a global threshold determined from the overall population FMR or FNMR. Alternatively, some works use yoking, where thresholds are computed using impostor pairs from the same demographic group, aiming to reflect intra-group calibration. In certain cases, thresholds are derived based solely on the FMR of a dominant group (e.g., White Caucasians as seen in FRVT [19]). Some works determine the threshold by fixing the maximum intra-group FMR to enhance robustness against demographic imbalance [95]. Since the threshold setting significantly affects all threshold-based fairness metrics, it is a crucial design choice that must be made with transparency and contextual awareness. As accuracy differentials become narrower with improvements in model architectures and training data, future fairness evaluations should include both threshold-based and threshold-agnostic metrics.
VI Bias Mitigation in Face Recognition Systems
Figure 4: Illustration of categories of bias mitigation methods in FR.
In an FR system, bias mitigation can be applied at different stages of the recognition pipeline, providing a structured approach to addressing demographic disparities. These stages align with general categories of bias mitigation strategies commonly used in machine learning: pre-processing, in-processing, and post-processing [24, 9, 111, 112].
Pre-processing methods aim to address fairness at the level of training data by modifying or augmenting it to reduce discriminatory patterns or imbalances before the data is used for training. This approach is particularly relevant when representational disparities in the data contribute to variations in FR performance across demographic groups. The in-processing methods focus on modifications to the FR model during training or fine-tuning phase, often by incorporating constraints or objectives that optimize fairness without significantly compromising recognition accuracy. Finally, post-processing techniques involve adjustments to the output of trained models to ensure fairness across demographic groups. These techniques modify the results, such as classification scores or decision thresholds, to achieve equitable performance without altering the underlying model. Figure 4 summarizes the categorization of bias mitigation methods– that correlates with the corresponding stages of FR pipeline. These methods are generally aimed at minimizing accuracy disparities across demographic groups while maintaining overall recognition performance. However, empirical evidence from most studies indicates that efforts to reduce such disparities often lead to a reduction in overall accuracy. Moreover, accuracy is often not the most meaningful metric in deployment settings, where systems are evaluated at specific operating points defined by fixing a particular error rate, such as FMR, FNMR, or TMR. As a result, bias mitigation techniques should as well consider application-relevant metrics evaluated at the corresponding operational thresholds.
VI-A Pre-Processing Methods
Pre-processing methods, also known as data-based methods, focus on modifying the biometric samples before feeding into the FR system. These techniques aim to normalize the characteristics of the data to make them robust for subsequent feature extraction. Most of the methods in this category can also be regarded as special case of data augmentation, specifically designed to reduce demographic disparities in the training data.
In [27], Klare et al. showed that FR algorithms were performing worse for female, Black, and younger individuals. To address this concern, they proposed two mitigation strategies based on selection of training data. First, training models on specific demographic cohorts to enhance recognition for those groups, and second, implementing a dynamic face matcher selection approach, where different algorithms, each trained on distinct demographic groups, would be chosen based on probe information. A similar approach was followed by Deb et al. in the context of longitudinal study of FR [113]. Using the data of more than 900 children captured over time, they demonstrated that recognition accuracy decreases as the time gap between image captures grows. To address this, they fine-tuned FaceNet [114] on a separate child face dataset, showing improved accuracy. The authors advocated the need for tailored training and evaluation of FR systems for different age groups. Lu et al. [41] investigated the influence of covariates–such as age, gender, pose, and skin tone– toward the performance of face verification. Their findings showed that using gender information to curate training data improves performance, particularly at low FMRs.
In [115], Kortylewski et al. demonstrated the effectiveness of using synthetic face images to mitigate the accuracy differentials arising from real-world datasets. By utilizing synthetic data for pre-training the FR model, they showed that the negative effects of dataset bias, particularly with regard to pose variations, could be significantly reduced. Their experiments revealed that pre-training the CNNs with synthetic data, the need for real-world data can be reduced by up to 75%. Although this study did not directly address demographic fairness, we include it in the review as it highlights the potential of synthetic data in improving the generalization performance of FR systems and reducing (non-demographic) dataset bias. Yucer et al. [116] proposed a GAN-based adversarial image-to-image augmentation method to generate racially balanced training data at subject level while preserving identity features. Evaluated across multiple loss functions, their approach reduced performance variance across racial groups and improved recognition accuracy for minority subjects. Recently, Kotwal and Marcel proposed an Image-to-Image transformation module called Demographic Fairness Transformer (DeFT), which enhances image representations before passing them to pretrained CNNs [117]. The DeFT uses multi-head encoders and soft-attention mechanisms to selectively enhance images based on inferred demographic information. The demographic labels of race or ethnicity are often non-discrete– but this concern has rarely been addressed. In [117], they replaced hard labels with probabilistic weights which are implicitly inferred at run-time. Their experiments show that DeFT reduces variations in FR performance across demographic groups and, thus improves model fairness, with some models also achieving slightly better accuracy compared to baseline systems.
VI-B In-Processing Methods
This category of works, also known as model-based methods, are applied during the feature extraction stage by modifying the weights of the FR model. The goal is to learn weights that generate features or embeddings that are less sensitive to demographic differences.
Amini et al. [118] proposed a debiasing algorithm that adjusts the sampling probabilities of data points in large datasets to reduce hidden biases. In the face detection use-case, their algorithm led to a decrease in race and gender bias while improving classification accuracy. To our knowledge, similar approaches have not been tested for recognition or verification applications. In [83], Wang and Deng introduced a reinforcement learning-based race balance network (RL-RBN) where they applied adaptive margins through deep Q-learning. Their method aimed to reduce the skewness of feature scatter between racial groups, leading to more balanced performance across different demographics. As a part of this work, they also released two datasets— BUPT-GlobalFace and BUPT-BalancedFace datasets— that were specifically designed to study racial bias in FR systems.
Wang et al. explored the issue of highly-skewed class distributions in FR datasets [119]. They proposed Large Margin Feature Augmentation (LMFA) and Transferable Domain Normalization (TDN) as methods to balance class distributions by augmenting and normalizing the feature space. These methods were shown to enhance the performance of underlying FR models by mitigating issues arising from class imbalance, which often correlates with demographic bias in under-represented groups. In [120], Yin et al. introduced a center-based feature transfer framework to address the under-representation of certain demographic groups in FR datasets. By transferring feature distributions from well-represented classes to under-represented ones, they augmented the feature space for these groups, reducing bias and improving recognition performance for under-represented subjects.
Gong et al. proposed a group-adaptive classifier (GAC) that uses adaptive convolution kernels and attention mechanisms tailored to different demographic groups [121]. By applying kernel masks and attention maps specific to each group, their method activates facial regions that are more discriminative for each demographic, thereby improving recognition accuracy and fairness across demographic groups. Another approach by the same authors [105] introduced an adversarial network, called DebFace, which utilizes a multi-task learning framework to simultaneously learn identity and demographic attributes. Their method employed adversarial training to disentangle identity features from demographic attributes such as gender, age, and race to effectively reducing differential in the recognition process. Their experiments demonstrated that, for DebFace, not only recognition but also demographic attribute estimation tasks were less biased. Huang et al. [122] introduced GABN, a de-biasing network that combines gradient attention maps (GAM) with adversarial learning to mitigate racial bias in face recognition. The method enforces consistency of GAMs across racial groups and uses GAM-guided sensitive region erasure to enhance feature learning for darker-skinned individuals. Their experiments showed that GABN was effective at reducing the performance gap between darker-skinned subjects and Caucasians. In [123], authors proposed unsupervised data partitioning, in iterative manner, followed by Invariant Feature Regularization, to improve generalization across demographic groups. The unsupervised partitioning was chosen as an inexpensive alternative to annotating training data with multiple (confounding) attributes.
TABLE V: Summary of recent works on mitigation of demographic bias in face recognition.
Reference Year Type of Method Test Dataset Summary
Klare et al. [27] 2012 Data-Processing PCSO Training models on specific demographic cohorts
Deb et al. [113] 2018 Data-Processing CLF Finetuning models on specific cohort (age, in this case)
Lu et al. [41] 2019 Data-Processing IJB-B, IJB-C Curation of training data
Kortylewski et al. [115] 2019 Data-Processing Multi-PIE, LFW, IJB-A Synthetic data pretraining followed by real data fine-tuning
Yucer et al. [116] 2020 Data-Processing RFW Adversarial image-to-image augmentation
Kotwal & Marcel [117] 2024 Data-Processing RFW Demographic-dependent transformation of input image
Amini et al. [118] 2019 In-Processing PPB Sampling data probabilities for face detection
Alasadi et al. [124] 2019 In-Processing CelebA, UMD Faces Introduces an adversarial learning framework using demographic classifier
Liang et al. [125] 2019 In-Processing CelebA Two-stage adversarial bias mitigation through disentangled representations and additive adversarial learning
Wang et al. [119] 2019 In-Processing RFW Large-margin feature augmentation to balance class distributions
Yin et al. [120] 2019 In-Processing LFW, IJB-A, MS-Celeb-1M Feature transfer to enhance under-represented groups
Wang & Deng [83] 2020 In-Processing RFW Reinforcement learning-based race balance network
Gong et al. [105] 2020 In-Processing RFW Adversarial debiasing with identity and demographic classifiers
Georgopoulos et al. [82] 2020 In-Processing KANFace Method to disentangle task-relevant features from bias terms in embeddings
Li et al. [126] 2021 In-Processing RFW, BFW Progressive cross-transformer to remove race-induced identity-unrelated components
Gong et al. [121] 2021 In-Processing RFW Group-adaptive training with adaptive convolution kernels and attention mechanisms
Xu et al.  [127] 2021 In-Processing RFW, BFW Additional penalty for disparities in FPR across instances
Wang et al. [128] 2021 In-Processing BUPT-Balancedface, Globalface Meta Balanced Network (MBN): a bilevel optimization framework that learns adaptive margin values per skin-tone group
Park et al. [129] 2022 In-Processing CelebA, UTK-Face Contrastive setups to enhance intra-class similarity and diminish similarity between negative samples
Serna et al. [130] 2022 In-Processing DiveFace, RFW, BUPT-Balancedface A triplet-based discrimination-aware loss, applicable to pre-trained FR models
Ma et al. [123] 2023 In-Processing MFR, CelebA, RFW Unsupervised data partitioning and Invariant Feature Regularization
Zhang et al. [131] 2023 In-Processing CelebA, UTK-Face Generating contrastive sample pairs with visual similarity and unsupervised feature reweighting
Huang et al. [122] 2023 In-Processing RFW Enforces consistent gradient attention across demographic groups and follows attention-guided erasure to enhance feature learning
Wang et al. [132] 2023 In-Processing RFW Minimizes performance gaps across identities using MixFair Adapter- without relying on sensitive attribute labels
Dooley et al.  [46] 2023 In-processing VGGFace2, CelebA Uses Neural Architecture Search and Hyperparameter Optimization to discover fair FR architectures
Kotwal & Marcel [107] 2024 In-Processing VGGFace2, MORPH, RFW Regularization constraints based on score calibrations for demographic groups
Li et al.  [133] 2025 In-Processing RFW, BFW, NFW Customized instance margins to preserve consistency in FPR and TPR during the label classification in training
Continued on the next page.
TABLE VI: (Continued)
Reference Year Type of Method Test Dataset Summary
Michalski et al. [134] 2018 Post-Processing Private Dynamic thresholds based on age differences
Srinivas et al. [135] 2019 Post-Processing ITWCC-D1 Score fusion and ensemble strategies to address age-related bias
Robinson et al. [86] 2020 Post-Processing BFW Demographic-specific thresholds
Terhörst et al. [106] 2020 Post-Processing Color-Feret, LFW Fairness-driven NN classifier
Morales et al. [84] 2020 Post-Processing DiveFace, CelebA Disentangling sensitive attributes from features for privacy preservation
Terhörst et al. [136] 2020 Post-Processing Color-Feret, MORPH Fair score normalization to mitigate demographic bias
Dhar et al. [137] 2021 Post-Processing IJB-C Privacy-preserving disentanglement of features via adversarial training
Salvador et al. [103] 2021 Post-Processing RFW, BFW Conditional calibration of clustered embeddings
Conti et al. [95] 2022 Post-Processing LFW Employing shallow MLP on the FR embeddings
Liu et al. [138] 2022 Post-Processing RFW Threshold Consistency Penalty (TCP) to enforce consistent decision boundaries across multiple datasets
Robinson et al. [85] 2023 Post-Processing BFW Privacy preserving domain adaptation framework reproject embeddings
Linghu et al. [139] 2024 Post-Processing VGGFace2, RFW Integrating demographic information in Z/T score normalization
Conti & Clémençon [140] 2024 Post-processing RFW Centroid-based loss aligns subgroup performance to a reference group via a lightweight fairness module
Another approach for domain-specific bias mitigation using disentangled representation learning was proposed by Liang et al. [125]. They introduced a two-stage method combining modules for disentangled representation learning with additive adversarial learning (AAL). While this work does not directly address accuracy differential, it provides useful insights into how domain-specific biases can be mitigated by learning disentangled representations. The effectiveness of this method in reducing bias across various domains suggests its potential applicability in the context of demographic fairness in FR. In [124], the authors introduced an adversarial deep learning framework for face matching aimed at reducing demographic bias. The model incorporated a demographic classifier adversary during training to encourage the learned feature representations to be invariant to the demographic attribute (gender, in their work). Their results showed that the adversarial setup reduced disparities in matching performance across demographic groups while maintaining accuracy comparable to a standard baseline.
In [126], a Progressive Cross Transformer (PCT) was proposed to mitigate racial bias by decoupling face representations into identity-related and race-induced components. Using dual cross-transformers, the PCT refines identity features and suppresses racial noise, demonstrating lower racial differentials without compromising recognition accuracy. In [82], the authors proposed a feature decomposition-based framework that separates task-relevant features from bias-related components in deep embeddings. In addition to FR, they applied this framework to age estimation and gender classification, demonstrating that certain demographic attributes can be disentangled without significantly degrading task performance.
A bias mitigation strategy leveraging Neural Architecture Search (NAS) and Hyperparameter Optimization (HPO) for fair FR models was proposed in [46]. By constructing a tailored search space, their authors synthesized architectures that generalize across datasets and protected attributes, such as gender. The models resulted from this work demonstrated reduced linear separability of demographic attributes.
The concept of score normalization was incorporated as a regularization term into the training objective enabling simultaneous optimization of recognition accuracy and demographic fairness [107]. This was facilitated by constraining the output scores of mated and non-mated pairs to adhere to a pre-defined distribution, and followed by minimizing differences in score distributions across demographic groups. During inference, the overall pipeline did not require modifications as the FR CNN architecture remained unaltered while only the weights were fine-tuned to the new objective.
A couple of works [127, 133] have proposed customized instance margins aimed at preserving consistent error rates– either false positives or false positives combined with true positives. These methods modify the Softmax-based loss by additional penalty terms derived from disparities in the concerned error rates. Wang et al. [128] proposed Meta Balanced Network (MBN), a meta-learning framework that dynamically adjusted margin values for different skin tone groups using gradients from a small, unbiased validation set. By framing margin optimization as an outer loop and model training as an inner loop, MBN learnt adaptive margins that improve fairness across groups. The bilevel optimization enabled their model to reduce bias while maintaining performance. Serna et al. [130] proposed Sensitive Loss, a discrimination-aware extension of the triplet loss function that incorporated a sensitive triplet mining strategy. Designed as an add-on to pre-trained FR models, this method enabled fairness improvements without retraining entire architectures. Wang et al. [132] propose MixFairFace, a framework designed to reduce identity bias– performance inconsistencies across individual identities– rather than relying on demographic group labels. The core component, MixFair Adapter, uses a sample-mixing strategy to estimate and minimize identity-level bias during training. This label-free approach enables fairness enhancement without requiring sensitive attribute annotations like race or gender.
Finally, we discuss a couple of works dealing with fairness in facial attribute recognition. Both of these are based on contrastive learning– which can be useful mechanism to address the demographic fairness in FR as well. Park et al. addressed fairness issues in attribute classification using a contrastive learning framework [129]. They constructed a Fair Supervised Contrastive Loss (FSCL) which reduces disparities by normalizing intra-group compactness and inter-group separability, penalizing sensitive attribute information in representations. Although predominantly an in-processing method, the authors also incorporate a loss function to address the imbalance in training data where majority groups are constrained to have a better intra- group compactness and inter-class separability compared to the under-represented ones. Similarly, Zhang et al. proposed Fairness-aware Contrastive Learning (FairCL) for unsupervised representation learning and demonstrated the use-case of facial attribute recognition [131]. In addition to fair contrastive learning of feature representations, they also attempted to address the dataset bias by specifically generating contrastive sample pairs that share the same visual information apart from sensitive attributes. They also suggested unsupervised feature reweighting to strike balance between the utility and fairness of learned representations.
VI-C Post-Processing Methods
These methods are applied after the feature extraction or matching stages to adjust decision scores and mitigate bias. Post-processing approaches generally fall into two categories: those that modify the features (embeddings) and those that operate directly on the similarity scores or their distributions. While these methods are less commonly used compared to pre- and in-processing techniques, they can still play a role in ensuring equitable outcomes. Additionally, post-processing methods are generally easy to integrate into existing FR pipelines. Their advantages include avoiding the need to retrain large models from scratch, enabling fairness adaptation using a smaller dataset, faster training times, and the ability to retain the performance of (a strong) pre-trained backbone. However, since the feature extractor remains fixed, these methods may offer limited capacity to correct upstream bias or improve learned representations.
Morales et al. [84] introduced SensitiveNets, a framework that enforces the removal of sensitive attribute information from learned feature representations. By disentangling sensitive attributes from task-relevant features, their method ensures that downstream agents cannot exploit protected information, thereby enhancing both privacy and fairness. Salvador et al. proposed FairCal: a post-processing method aimed at improving fairness in face verification without retraining or requiring sensitive attribute labels [103]. They clustered image embeddings (obtained from frozen FR network) in an unsupervised manner and applied conditional calibration maps per cluster. They reported reduced false matches while improving overall accuracy. In [95], Conti et al. employed a shallow MLP on top of the frozen FR model as a post-processing strategy to mitigate gender bias in FR. They enhanced the representations of underperforming subgroups using a Fair von Mises-Fisher loss with class-specific variance control. The work in [140] introduced a post-processing method that improves fairness in pretrained (frozen) FR models using a centroid-based loss called Centroid Fairness. A lightweight Fairness Module was trained to align subgroup performance with that of a reference group, without modifying the original model.
Feature descriptors, or embeddings, generated by FR networks inherently encode sensitive identity-related attributes, such as gender and skin-tone; thereby rendering them vulnerable to privacy risks. Although the preservation of privacy lies outside the scope of this review, it is closely linked to concerns of fairness, given its role in exacerbating performance disparities across demographic groups. Research on disentangling sensitive attributes from identity-related representations remains relatively sparse. In [141], Dhar et al. proposed the adversarial gender de-biasing algorithm, which aims to reduce gender-specific information from embeddings extracted by pre-trained (frozen) FR networks. This method employs adversarial training to simultaneously suppress gender information and maintain identity recognition performance. Building on this approach, the Protected Attribute Suppression System (PASS) [137] advances the concept by introducing a descriptor-level adversarial framework to mitigate accuracy differentials. PASS operates directly on pre-trained embeddings and employs a novel discriminator-based technique to suppress attributes such as gender and skin tone without necessitating end-to-end retraining. Additionally, Robinson et al. [85] highlighted that applying a single decision threshold across demographic subgroups results in disparate performance, diverging from global evaluation metrics. To address this, they proposed a domain adaptation framework that re-projects embeddings into a feature space where demographic attributes are suppressed, while identity-relevant information is retained.
Michalski et al. investigated the impact of age variation on FR, particularly for children [134]. They showed that dynamic thresholding improves performance. To address age-related accuracy differences, they adjusted thresholds based on age differences as opposed to a fixed threshold. In another work on age-related bias, Srinivas et al. experimented with score-level fusion strategies to improve recognition accuracy for the children (age as demographic) [135]. They considered six fusion schemes that combined different score-normalization techniques and fusion rules. For normalization, they considered $\mathbb{Z}$-norm and $\min$-$\max$ strategies; while fusion was conducted using min, max, or sum rules. Robinson et al. [86] showed that applying a single threshold across different demographic groups leads to significant variations in the FMR. They addressed this issue by using per-subgroup thresholds to balance the FMRs across ethnic and gender groups, improving both recognition fairness and performance.
A typical FR pipeline employs similarity functions to obtain a matching score. Terhörst et al. replaced the conventional similarity function by a fairness-driven neural network classifier [106]. By adding a penalization term in the loss function, their method was able to equalizes score distributions across ethnic groups, reducing intra-ethnic bias while maintaining high recognition performance. In another work [136], Terhörst et al. introduced an unsupervised fair score normalization method based on individual fairness principles, which treats similar individuals similarly. During training, they partitioned the identities in finite number of groups using K-means clustering on face embeddings. At inference, they computed the cluster-specific thresholds for both samples contributing to the score, and these threshold were combined with a global threshold to yield normalized scores. In a recent work, Linghu et al. extended traditional score normalization methods, such as $\mathbb{Z}$ and $\mathbb{T}$ normalization, by incorporating demographic information to enhance fairness in FR systems [139]. Furthermore, they evaluated three cohort-based approaches based on imposter scores, Platt scaling, and bi-modal cumulative distribution functions (CDF). Their findings demonstrated that the proposed method improved fairness across both race and gender demographic groups, particularly at low FMRs.
Liu et al. [138] argued that the existing FR evaluation protocols, which rely on dataset-specific thresholds, are practical, and they introduced the One-Threshold-for-All (OTA) protocol that employs a single calibration threshold across all datasets. To optimize performance under the OTA setting, they also proposed the Threshold Consistency Penalty (TCP), which enforces consistent decision boundaries across domains through implicit domain division followed by calibration and estimation of domain thresholds.
The works on mitigation are summarized in Table V. For most of the state-of-the-art works aimed towards mitigating accuracy differentials, the FR models have been trained on massive datasets, fine-tuned primarily for accuracy, and are potentially operating near their capacity limits. When a new objective, such as fairness, is introduced, it competes with the existing accuracy objective, often resulting in performance degradation. This effect is commonly known as the performance-fairness trade-off, where improvements in fairness are often accompanied by reduced recognition accuracy.
Some recent methods have demonstrated the potential to improve fairness metrics without sacrificing overall accuracy. These methods, however, often involve architectural modifications, such as the addition of auxiliary modules. It is important to note that in end-to-end FR pipelines, the total number of model parameters can vary significantly depending on the chosen bias mitigation strategy. Some in-processing methods aim to improve fairness by regularizing or fine-tuning the existing FR network, without adding any additional modules. In contrast, other approaches may integrate substantial additional components, potentially increasing the network depth and capacity. While such designs may yield improved recognition performance along with reduced bias, they differ significantly in architecture and complexity from the baseline models. In addition to structural differences, variations in training data used for both– bias mitigation and FR backbone– play important role. As a result, performance comparisons may not be directly equivalent, and should be made with transparency and careful consideration.
VII Future Directions
In recent years, with advanced architectures and increase in the number of layers and parameters, FR models have gained a substantial improvement in their capacity to learn complex facial representations. These deeper architectures, often comprising over a hundred layers with millions of parameters, have significantly enhanced the ability of FR systems to generalize across challenging scenarios, resulting in higher overall recognition performance (as well as improved fairness). Additionally, the availability of larger, more diverse datasets has contributed to better learning outcomes. These datasets, which incorporate substantial variation in demographics and covariates such as pose, illumination, and expression (PIE), have facilitated measurable progress in improving both accuracy and fairness FR.
However, despite these advancements, several key challenges in this area continue to exist. Most of the existing efforts primarily center on extremely deep models, which demand extensive computational resources and memory footprint for both training and deployment. This emphasis on high-capacity architectures does not adequately address quality issues in data or labels nor does it cater to the requirements of resource-constrained environments. Thus, while the combination of deeper models and diverse data has been pivotal, future research must explore avenues to address residual biases and expand fairness to a broader range of applications.
In this section, we examine some of the emerging challenges associated with demographic fairness. These challenges highlight the need for ongoing research to adapt bias mitigation strategies to align with the advancements in FR applications.
Lightweight Models: The lightweight FR models, often used in handheld devices and resource-constrained environments, encounter significant challenges concerning demographic fairness. These systems, crucial for privacy-sensitive applications, often inherit limitations in capacity and architecture, leading to non-equitable performance across demographic groups. Performance disparities in lightweight models has garnered limited research attention, despite their widespread deployment in mobile and IoT devices with varying sensor qualities. Techniques like knowledge distillation (KD) and pruning, while essential for model compression, introduce or amplify bias. For instance, Liu et al. highlighted that KD inherits biases from larger teacher models [142], while pruning has been shown to disproportionately impacts underrepresented groups [143, 144]. Incorporating fairness-aware techniques is crucial for mitigating these issues. Lin et al. [145] introduced FairGRAPE, a pruning method that evaluates network connections with demographic considerations, reducing performance disparities. Caldeira et al. [146] proposed a multi-specialized teacher framework where each teacher model is trained on a specific ethnicity and collectively distills knowledge into a student model. Achieving demographic fairness in lightweight models requires targeted compression strategies and ethical evaluations of demographic-specific impacts. By integrating fairness principles into compression techniques, lightweight FR systems can achieve more equitable outcomes while maintaining efficiency and accuracy.
Quantization of Models and Data: Quantization, a model compression technique, retains the original architecture (as opposed to KD and pruning– which often modify the structure) while reducing parameter precision, producing smaller and faster models. However, it can lead to higher disparities across demographic groups by prioritizing global performance over the accurate classification of under-represented groups. Such performance disparities underscore the need for rigorous fairness evaluation across demographic subgroups when deploying compressed models. Quantization converts floating-point (FP) models into lower-precision formats like 8-bit, balancing efficiency and accuracy [147]. It typically consists of two approaches: post-training quantization (PTQ) and quantization-aware training (QAT). Some studies, such as Stoychev et al. demonstrated that 8-bit PTQ maintained fairness and accuracy in gender bias for face expression recognition [147]. However, reducing precision to 6 bits significantly degraded fairness, indicating a trade-off between compression and bias mitigation. Although similar investigations for FR systems remain limited, these findings highlight the need to evaluate and ensure demographic fairness in quantized models.
While the majority of research in this area focuses on quantized models, a number of recent studies have investigated the implications of lossy image compression on fairness in FR. The work in [148] explored how lossy JPEG compression influences recognition accuracy across demographic groups, more specifically on racial phenotypes. Their findings indicate that compression disproportionately degrades performance for certain traits, such as darker skin-tones and wider facial features. They also demonstrated that disabling chroma sub-sampling can improve FMRs across several racial phenotypes. Similarly, Qiu et al. [149] examined the effects of five neural image compression methods and found that certain phenotypic features–such as skin-tone and hair-type–are more susceptible to degradation at lower bit rates. Notably, at extremely low bit rates, the amplification of bias was particularly pronounced for individuals belonging to African race.
Low Resolution: The research on demographic fairness in FR has predominantly focused on high-resolution images, often overlooking the challenges posed by low-resolution images typically captured by surveillance cameras or from significant distances. One of the primary impediment to research this issue is the scarcity of low-resolution datasets that include demographic attributes. Consequently, demographic disparities in low-resolution FR remain under-explored, despite their importance in real-world applications. A recent work from Atzori et al. attempted to address this gap by designing a novel framework to investigate performance differences across demographic groups in low-resolution FR [150]. They trained state-of-the-art FR models on various combinations of high- and low-resolution images. Testing on degraded images from five datasets revealed significant disparities across gender and ethnic groups, underscoring the need for timely interventions in low-resolution FR. It may be noted that their approach involved use of a generative model to convert high-resolution face images into realistic low-resolution counterparts. In [51, 66], the impact of blur and image resolution were investigated in the context of 1:$N$ identification, revealing that increased blurring leads to elevated FPIRs. These studies also noted that the impact of blur varies across demographic attributes, particularly race and gender. The quality degradations of blur and resolution in these studies were simulated using Gaussian blurring and bicubic interpolation, respectively. The importance of low-resolution FR is evident in programs like BRIAR [^3] BRIAR Programme by IARPA., which aim to enhance recognition technologies for challenging scenarios, such as long-distance identification and low-quality image acquisition. To tackle demographic fairness in low-resolution FR, there is a need to develop both datasets and models tailored to these unique use-cases.
Datasets for Training and Evaluation: While demographically balanced training datasets alone may not fully eliminate performance differences in FR, they play a critical role in reducing such disparities. Achieving fairness requires not only demographic diversity (e.g., across race, gender, and age) but also balance with respect to non-demographic or soft attributes such as hairstyle, facial hair, makeup, and occlusion. Therefore, large-scale datasets that are balanced across both demographic and non-demographic dimensions are essential for developing FR models that are both fair and accurate. However, acquiring such datasets is increasingly challenging due to cost and ethical and privacy concerns surrounding biometric data collection. The commercial sector, in particular, faces difficulties as most available datasets are collected locally from consenting individuals—which are often in limited size and demographic representation. These constraints necessitate innovative approaches to address demographic fairness using smaller or synthetic datasets. When it comes to evaluation and benchmarking, many existing datasets are significantly imbalanced– typically over-representing a single demographic group, such as Caucasians. Alongside developing more inclusive training datasets, there is a pressing need to construct evaluation datasets that more accurately reflect demographic diversity.
Access to datasets annotated with multiple attributes–both demographic and non-demographic–will prove to be a valuable resource for advancing fairness in biometric recognition. Such datasets enable deeper exploration of the underlying causes of performance differences across demographic groups by facilitating the analysis of correlations and potential causal relationships among factors that influence recognition performance. Moreover, they provide a foundation for designing effective mitigation strategies. When used for assessment, multi-attribute datasets can help controlling or isolating individual factors, enabling systematic benchmarking of their specific impact on model performance.
The use of synthetic data in FR has recently gained traction as a potential solution to privacy and data-sharing concerns. Competitions like the FRSyn series have encouraged advancements in synthetic data usage [109, 110, 108]. Despite these efforts, FR models trained exclusively on synthetic datasets continue to underperform compared to those trained on real datasets of similar size [108, 151]. This gap is evident in both recognition accuracy (measured by metrics like FMR and FNMR) and demographic fairness (assessed by standard deviation of performance across groups). The analysis of synthetic datasets by Huber et al. revealed that demographic bias might worsen compared to the (real) training dataset [152]. Enhancing the quality and utility of synthetic datasets, beyond the aspect of fairness, remains an open problem, requiring further exploration.
Causes of Performance Disparities: Recent research indicates that accuracy disparities across demographic groups in FR systems are not solely attributable to inherent physical or biological characteristics. Instead, these disparities are frequently influenced by non-demographic factors– such as hairstyle, cosmetic use, or occlusion patterns– that are shaped by cultural norms, social behaviors, or individual preferences. Despite their relevance, such factors have historically received limited attention, with only a few recent studies beginning to reveal their significant impact on algorithmic performance.
A comprehensive understanding of the origins of demographic accuracy differentials requires disentangling the complex interplay between demographic attributes and confounding behavioral or environmental variables. This necessitates the development of training and evaluation datasets that are not only demographically diverse but also annotated for a broader set of non-demographic characteristics potentially correlated with recognition performance.
To systematically analyze these influences, it is essential to design datasets that enable controlled isolation– analogous to evaluating a partial derivative– of individual or combined factors. This would allow researchers to measure the effect of specific attributes or conditions on the performance of a FR model or pipeline.
Advancing our understanding of these causal relationships is critical not only for identifying the root causes of observed biases but also for facilitating systematic and reproducible evaluation. Ultimately, understanding these underlying causes is a prerequisite for developing effective mitigation strategies and building FR systems that are both fair and accurate.
Use-Cases of Remote Checking: Last few years have witnessed tremendous surge in online activities: financial transactions, banking, user-onboarding, etc. These activities have driven widespread adoption of remote identity verification (RIdV) technologies. These systems authenticate individuals by comparing real-time images or selfies, captured via smart devices, against official identity documents, such as work permits or driver’s licenses. Such solutions are integral to online Know Your Customer (KYC) processes, which are now standard for banks and financial institutions. While RIdV systems enhance convenience and scalability, it is essential to ensure their fairness across demographic groups as they become more prevalent.
Recognizing the increasing reliance on remote verification, the MdTF and DHS S&T introduced the Remote Identity Validation Technology Demonstration (RIVTD) initiative [^4] Remote Identity Validation Technology Demonstration (RIVTD). In addition to security, accuracy, and liveness detection requirements, this program also places particular emphasis on ensuring demographic fairness in such technologies. A recent study by Fatima et al. [153] investigated demographic fairness in RIdV technologies using statistical methods to analyze performance disparities. Their analysis of five commercial RIdV systems revealed that only two achieved equitable outcomes across demographic groups. Notably, higher FNMRs were observed among African cohorts and individuals with darker skin-tones. Such findings highlight the necessity of evaluating RIdV technologies across demographic groups to ensure equitable and fair performance.
Complex Bias Factors (Intersectionality): The majority of research on mitigation of accuracy differentials in FR, as discussed in Sec VI has focused on single demographic attributes, such as race, age, or gender. However, several studies have identified that combination or intersection of various demographic factors causes (or amplifies) bias in FR models (cf. Sec III), while few works, such as [69, 52] have addressed the issue of disparities caused due to multiple or combined attributes. Existing bias mitigation techniques typically target one demographic attribute at a time, achieving measurable improvements in fairness for that specific attribute. However, in most cases, it remains unclear whether such processing inadvertently introduces imbalances in other demographic attributes. For instance, enhancing fairness for gender-related bias may increase disparities linked to ethnicity and vice-versa. This highlights the need for systematic evaluations of the intersectionality of demographic factors, such as race and gender combined. Consequently, developing mitigation methods capable of addressing multiple demographic attributes simultaneously remains an open challenge.
Noisy Labels: The assignment of demographic attributes such as race, ethnicity, and skin-tone in FR datasets typically involves discrete labeling into finite categories. Some attributes, such as race, are often self-reported. In many cases, race and ethnicity annotations may be derived from automatic classifiers or manual efforts. Automatic classifiers, predominantly based on deep learning, are likely to be susceptible to bias too; while manual annotations are prone to human judgment errors. Similarly, skin-tone is frequently categorized using scales like Fitzpatrick’s, which discretizes it into specific values, ignoring its continuous spectrum. This reliance on discrete labels introduces noise into the training data, as samples near category boundaries are often inaccurately labeled. Recent work [117], addressed this issue by using probabilistic weights (soft labels) for demographic information instead of utilizing rigid (categorical) labels. However, most existing methods overlook the issue of errors in training data.
Noisy labels in training datasets pose a significant challenge, especially considering massive scale of FR datasets, often in few hundreds of thousands of images. Manually verifying or curating such datasets is labor-intensive, impractical, and still prone to errors. Furthermore, removing samples with ambiguous labels can lead to reducing dataset diversity and robustness. Thus, developing robust mitigation strategies capable of handling noisy labels without compromising the effectiveness of training processes or dataset diversity is essential for improving fairness and accuracy in FR systems.
VIII Conclusion
In this work, we have systematically explored the issue of performance differences in FR across demographic groups, or bias through different yet interrelated sections: causes, datasets, assessment metrics, and mitigation strategies. We discussed key contributing factors such as training data imbalance, skin-tone variations, and image quality, as well as the growing recognition of non-demographic attributes. Our analysis of datasets emphasized the importance of demographic diversity and the need for multi-attribute annotations to enable deeper understanding of performance disparities. We also reviewed a range of fairness evaluation metrics and mitigation techniques across the FR pipeline.
As highlighted by several recent studies, non-demographic covariates (or soft-biometric factors) can have a significant influence on FR performance, including accuracy and error rates. Given that both training and evaluation datasets, including real-world deployments, are rarely controlled for demographic attributes in isolation, it becomes critical to reconsider how we interpret variations in FR outcomes across demographic groups. Many of these non-demographic factors are closely entangled with demographic attributes (e.g., hairstyles and makeup often vary by culture or gender). Consequently, performance disparities observed across groups may not be attributable to the demographic variable alone. Without precise isolation and identification of contributing factors, attributing lower FR performance to demographic bias may be misleading. Therefore, caution must be exercised before concluding that an FR system exhibits bias toward a particular demographic group.
Another important challenge to be noted here is the trade-off between fairness and accuracy. Since most FR models are already optimized for performance, adding fairness objectives often introduces conflicting constraints. Although some architectures sometimes improve both, these gains frequently rely on increased model capacity, rather than an inherent resolution of the fairness-accuracy trade-off. Recent studies have shown that gender-related performance differences may be explained by soft-biometric attributes rather than inherent biological differences, suggesting that some disparities are socio-cultural in nature. However, such understanding is still limited for attributes like race or skin-tone, largely due to the absence of datasets annotated with both demographic and non-demographic variables. Creating such datasets is essential but resource-intensive; synthetic data may help address this gap, though its effectiveness remains sub-par of real-world data. Improved metrics are also needed to quantify the impact of specific attributes on performance in a consistent and interpretable way. Additionally, most fairness evaluations are conducted on large-scale models, yet future deployments–especially on mobile and edge devices–will require lightweight models that may be more prone to differential treatment. Ensuring fairness under such constraints is critical for real-world adoption.
Together, these challenges reflect the evolving nature of FR technologies and underscore the need for innovative strategies to ensure equitable and reliable outcomes in real-world applications.
Acknowledgment
The work has been supported by the Hasler foundation (through the SAFER project) and the Swiss Center for Biometrics Research and Testing.